### v1

# MAVias: Mitigate any Visual Bias

Abstract
Mitigating biases in computer vision models is an essential step towards the trustworthiness of artificial intelligence models. Existing bias mitigation methods focus on a small set of predefined biases, limiting their applicability in visual datasets where multiple, possibly unknown biases exist. To address this limitation, we introduce MAVias, an open-set bias mitigation approach leveraging foundation models to discover spurious associations between visual attributes and target classes. MAVias first captures a wide variety of visual features in natural language via a foundation image tagging model, and then leverages a large language model to select those visual features defining the target class, resulting in a set of language-coded potential visual biases. We then translate this set of potential biases into vision-language embeddings and introduce an in-processing bias mitigation approach to prevent the model from encoding information related to them. Our experiments on diverse datasets, including CelebA, Waterbirds, ImageNet, and UrbanCars, show that MAVias effectively detects and mitigates a wide range of biases in visual recognition tasks outperforming current state-of-the-art.
Figure 1: MAVias identifies visual biases through foundational models that extract tags representing visual features and assess relevance to the target class. Then, MAVias encodes these features within the vision-language space and integrates them into a bias-aware framework to train a model that is invariant to such visual biases.
1 Introduction
Computer Vision (CV) progress has been largely driven by Deep Learning (DL) advances [9, 15] and large-scale datasets [6, 25, 21], enabling models to learn complex patterns and visual features with impressive accuracy. However, alongside this progress, concerns have emerged about biases embedded in these models [32, 29, 11, 38, 41, 44] – often stemming from unintended correlations present in the training data [27, 13, 30, 22]. The correlated attributes that are irrelevant, act as “shortcuts” and can significantly impact the model’s reliability and generalization [37, 12]. It is important to note that in the context of this paper, we define as visual bias any characteristic that does not contribute to defining the target class, which we refer to as “irrelevant”. To address this, several methodologies have been developed. Broadly, these fall into two categories: Bias Label-Aware (BLA) and Label-Unaware (BLU) methods. BLA methods leverage the annotations of the attributes introducing the biases to address them [19, 45, 42, 2]. BLU methods focus on extracting bias pseudo-labels in cases of extreme biases where a bias-proxy model (or bias-capturing classifier) can be trained using the task’s target labels that closely align with the bias labels [16, 31, 39, 28].
While effective in certain contexts, both BLA and BLU methods are limited in their applicability when there are multiple, complex, and possibly unknown biases. Common challenging scenarios include:
Unknown biases. In large-scale, general-purpose CV datasets, such as ImageNet, biases may be difficult to identify and largely remain unknown, as they vary widely between different classes and are not prominent enough to allow training of a bias proxy model. For instance, in the ImageNet9 example presented in Tab. 1, a sample labeled as a dog could introduce biases related to the background scene (e.g., armchair, couch, pillow, red), the color of the dog (e.g., black and white), or accessories like a neckband.
Potential biases beyond a predefined set. The CelebA example in Tab. 1, shows that beyond the hair color, additional biases may be present, such as clothing styles (e.g., business suit, tie).
Poor representation of predefined bias. In some cases, biases are reduced to single labels, such as “rural background” in the UrbanCars dataset. However, as shown in Tab. 1, more nuanced descriptors at the instance level (e.g., path, tree, wood, forest, hydrant, red, etc.) provide a richer context for modeling bias.
Existing BLA and BLU methods are not designed to mitigate such biases, leading to models that are biased and/or do not achieve their optimal generalization potential. To address this, Mitigate Any Visual bias (MAVias) introduces a flexible and scalable solution, capable of identifying and mitigating open-set biases in CV datasets. MAVias begins by extracting descriptive tags that capture various visual features, such as general-purpose objects, actions, scenes, and visual attributes. Note that a sufficiently large and comprehensive vocabulary can cover the vast majority of visual features present in CV datasets, enabling the method to be considered as operating in an open-set setting. Then, these tags are processed by a Large Language Model (LLM) to identify which ones are irrelevant to each of the target classes, leading to a rich set of language-encoded visual biases, text descriptions of visual characteristics that are irrelevant to the classification task at hand. We then translate these biases into vision-language embeddings, project them to the main model’s backbone space and then to the classification layer. This projection layer is trained simultaneously alongside the main model, and during training, the output logits are a linear combination of the logits of the main model and those of the projection layer that capture visual biases. This setup allows the main model to be exposed to the biased features – those representing irrelevant information to the target class – in a controlled way that leads to bias-invariant representations. Overall, MAVias provides an effective, end-to-end solution for identifying and mitigating biases in open-set scenarios. We evaluate the proposed method on several datasets involving single-attribute biases (CelebA, Waterbirds) as well as multi-attribute predefined biases (UrbanCars) demonstrating state-of-the-art performance. Furthermore, experiments were conducted on ImageNet9, involving unknown biases, where the suggested approach demonstrates significant gains (from 5.24% to 10.21%) in terms of accuracy compared to existing competitive approaches.
The main contributions of this paper are the following: (i) a framework for identifying instance-specific open-set potential visual biases in CV datasets, (ii) a learning strategy that exploits foundation models to learn bias-invariant representations, and (iii) an extensive evaluation study including comparative and ablation experiments across 4 thematically diverse datasets demonstrating the effectiveness and general applicability of the proposed method, which outperforms the current state-of-the-art.
2 Related Work
Bias identification
Several methods recently emerged that leverage text (such as captions, keywords, or tags) for bias detection, highlighting the potential of this approach in this domain. For instance, Say My Name (SaMyNa) [4] is an explainability method that tries to discover model biases through text-based pipelines. Similarly, the Bias-to-Text (B2T) [20] and Language-guided (Lg) [50] frameworks discover biases by extracting common keywords from the captions of misclassified images. These methods aim to identify the main source of bias (e.g., hair color in the CelebA example illustrated in Fig. 1) rather than discovering any potentially biased visual characteristic in an open-set setting. Furthermore, OpenBias [8] introduces a framework for detecting biases in text-to-image generative models by leveraging LLMs to propose potential biases from captions. Although operating in a different domain (text-to-image generation), similarly to MAVias  OpenBias acknowledges the need for discovering potential biases in an open-set setting. However, it solely relies on an LLM to generate specific bias lists for a given caption, raising concerns about whether the LLM’s interpretation of biases aligns with human judgment. In contrast, MAVias takes a more structured and systematic approach; it evaluates each descriptive tag by querying the LLM to determine whether it is directly relevant to the target class using a detailed prompt (refer to the Appendix for system prompt details).
Bias mitigation
Recent efforts to mitigate bias in CV have led to the development of various methodologies, including methods with direct access to the labels of attributes introducing bias (i.e., BLA methods) [19, 45, 42, 16, 2, 37, 35, 24] and methods that do not take advantage of such labels but instead rely on deriving pseudo-labels (i.e., BLU methods) [3, 1, 31, 39, 46]. For instance, Learning Not to Learn (LNL) [19] is a BLA method that discourages the model from predicting the attribute introducing bias, while Bias Contrastive-Bias Balance (BC-BB) [16] and FairKL [2] rely on the bias labels to enforce bias-neutral pairwise similarities between the samples using contrastive learning. Some methods have indirect access to the bias labels by utilizing bias-capturing classifiers trained on different datasets offering bias-related information explicitly [39, 40]. Finally, other methods infer pseudo-labels from the biased vanilla model to identify the biases [31, 1, 5]. While these methods have been effective in mitigating biases, they primarily depend on predefined bias labels or pseudo-labels derived from biased models. On the contrary, this work investigates open-set scenarios and introduces a flexible bias mitigation methodology that can discover and handle multiple, diverse biases without requiring a unified bias-labeling system across all the dataset’s samples.
3 Methodology
3.1 Problem Formulation
Let $\mathcal{D}=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$ be a dataset consisting of $N$ images, where $\mathbf{x}^{(i)}\in\mathcal{X}$ represents an input image, and $y^{(i)}\in\mathcal{Y}$ is the corresponding target label. The goal is to train a DL model $f_{\boldsymbol{\theta}}:\mathcal{X}\to\mathcal{Y}$, parameterized by $\boldsymbol{\theta}$, to predict the target label $y^{(i)}$ given an image $\mathbf{x}^{(i)}$, while mitigating a set of potential biases, $\mathcal{B}^{(i)}$, present in $\mathbf{x}^{(i)}$ that may lead to biased predictions. In this context the term “potential biases” refers to all visual attributes present in $\mathcal{X}$ that are irrelevant to the target class, such as elements of the background of images.
3.2 Methodology
Table 1: Examples of extracted tags for various datasets. Red color indicates the irrelevant tags ($\mathcal{B}^{(i)}$).
Dataset ImageNet9 Waterbirds UrbanCars CelebA
Target Class dog bird species car type gender
Bias Type unknown background background and objects hair colour
Sample [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Extracted Tags armchair, black, chair, couch, dog, neckband, pillow, red, sit, white, bamboo, bamboo forest, bird, blue, branch, green, hide, parrot, perch, sit, stand, stem, tree, yellow, car, path, forest, hydrant, lush, park, red, road, sedan, silver, SUV, tree, white, wood, black, business suit, dress, dress shirt, man, stand, stare, suit, tie, wear,
3.2.1 Language-driven Bias Modeling
First, MAVias utilizes an image tagging model to extract descriptive tags that represent the visual information present in an image. In many general-purpose CV datasets, biases consist of visual information that can be described in text. Formally, for each image $\mathbf{x}^{(i)}$, we derive a set of tags $\mathcal{T}^{(i)}=\{t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{m}\}$, where $m$ represents the total number of tags. These capture various visual attributes, including colors, objects, backgrounds, and other features that can either describe the target class or introduce bias.
The next step is to filter out tags that should not influence the decisions of a robust classifier. To achieve this, we leverage an LLM to identify which tags are irrelevant to the target class $y^{(i)}$. We denote this subset of potential biases as $\mathcal{B}^{(i)}\subseteq\mathcal{T}^{(i)}$. These encapsulate visual features that could lead to biased predictions if considered by the model. To ensure the LLM correctly identifies these irrelevant tags, we carefully design the prompt used for this task by providing precise instructions on what tag should be considered relevant (e.g., physical components, defining features, inherent characteristics, etc.) or not (e.g., background details, lighting, textures, other objects, etc.). Details on the prompt formulation process are provided in the Appendix. Table 1 shows several examples of $\mathcal{B}^{(i)}$ for samples belonging to different datasets.
For each set of irrelevant tags $\mathcal{B}^{(i)}$ associated with an image $\mathbf{x}^{(i)}$, we employ a vision-language model to generate a single embedding $\mathbf{e}^{(i)}\in\mathbb{R}^{d}$, where $d$ is the dimension of the embedding. This is produced using the prompt “a photo of $t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{k}$”, where $k\leq m$ is the number of irrelevant tags. This aggregate embedding captures the combined information from all irrelevant tags, providing an instance-level representation of the biased features that could affect the model’s behavior. Comparison with alternative embedding approaches is provided in Appendix.
Figure 2: Illustration of the proposed framework for mitigating any visual bias during model training. For inference, only the backbone and the classification layer are considered (i.e., $f_{\boldsymbol{\theta}}$).
3.2.2 Bias Mitigation
We define the main model $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})$, a DL classifier composed of the following: (i) a backbone $f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})$, which extracts feature representations $\mathbf{h}^{(i)}\in\mathbb{R}^{r}$, where $r$ is the feature vector size; (ii) a classification head $f_{\boldsymbol{\theta}_{c}}(\mathbf{x}^{(i)})$ outputting the logits $\mathbf{z}^{(i)}_{\text{main}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{h}^{(i)})\in\mathbb{R}^{p}$, where $p$ is the number of classes. Thus, the overall main model is expressed as:
$f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})=f_{\boldsymbol{\theta}_{c}}(f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})).$ (1)
In parallel, we introduce a projection layer $g_{\boldsymbol{\phi}}$, parameterized by $\boldsymbol{\phi}$, which takes the visual bias embeddings $\mathbf{e}^{(i)}$ as input and outputs embeddings $\mathbf{b}^{(i)}\in\mathbb{R}^{p}$. Note that $g_{\boldsymbol{\phi}}$ is employed to project $\mathbf{e}^{(i)}$ to the feature space of $\mathbf{h}^{(i)}$. Then the corresponding logits are derived through $\mathbf{z}^{(i)}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{b}^{(i)})\in\mathbb{R}^{p}$. The final logits $\mathbf{z}^{(i)}$ for each sample are the addition of the main model logits $\mathbf{z}^{(i)}_{\text{main}}$ and the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$:
$\mathbf{z}^{(i)}=\mathbf{z}^{(i)}_{\text{main}}+\mathbf{z}^{(i)}_{\text{tag}}.$ (2)
It is worth noting that bias often arises during DL model training because biased attributes in the training data are easier to learn and thus dominate gradient updates [34]. To counteract this, MAVias incorporates visual bias logits into the main model’s logits, ensuring that as the bias in a sample increases, its impact on gradient updates is reduced. The intuition behind this mechanism is that for bias-aligned samples, the value of $\mathbf{z}_{tag}$ is high, effectively reducing the magnitude of $\mathbf{z}_{main}$ and its contribution to the total logits $\mathbf{z}$. This leads to significantly reduced gradients for these samples. This is supported both empirically in Sec. 4.2 and theoretically in the Appendix. In other words, $\mathbf{z}_{tag}$ assists in decoupling the learning of biased features from the actual task at hand, which allows the main model to focus on the relevant features for the target prediction.
Furthermore, since both the main model and the projection layer are trained concurrently, it is essential to ensure the stability of the training process. To achieve this, we introduce a loss function that combines the classification loss with a regularization term. The classification loss ensures that the combined logits $\mathbf{z}^{(i)}$ accurately predict the target label $y^{(i)}$, while the regularization term aligns the magnitude of the main model’s logits $\mathbf{z}^{(i)}_{\text{main}}$ with those of the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$. By doing so, we prevent $g_{\boldsymbol{\phi}}$ from dominating or being overshadowed by $h_{\boldsymbol{\theta}}$.
Specifically, the total loss is computed as:
$\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})+\alpha\cdot\mathcal{L}_{reg}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$
where: (i) $\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})$ is the classification loss (e.g., cross-entropy loss) between the final logits $\mathbf{z}^{(i)}$ and the ground truth label $y^{(i)}$; (ii) $\mathcal{L}_{reg}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$ is the regularization term for the norm of the logits, calculated as:
$\mathcal{L}^{(i)}_{\text{reg}}=\frac{1}{2}\left\|\|\mathbf{z}^{(i)}_{\text{main}}\|-\lambda\cdot\|\mathbf{z}^{(i)}_{\text{tag}}\|\right\|^{2}$ (3)
[…]
4 Experiments
4.1 Datasets and Evaluation Protocol
We employ Biased-CelebA [16], Waterbirds [37], UrbanCars [24], and Imagenet9 [47]. Biased-CelebA is a subset of the well-known CelebA dataset consisting of facial images annotated with 40 binary attributes. In this subset, gender is treated as the target attribute, while the BlondHair attribute introduces bias by enforcing a 90% correlation between the two. Waterbirds is a dataset that features a co-occurrence of 95% between waterbirds (or landbirds) and aquatic environments (or terrestrial environments) as background. UrbanCars is an artificially generated dataset featuring a 95% co-occurrence between car body types and background and/or certain objects relevant to urban or rural regions. ImageNet9 consists of 9 coarse ImageNet classes (dog, bird, vehicle, reptile, carnivore, insect, instrument, primate, and fish).
Regarding the evaluation, we follow the protocols suggested by the dataset providers and previous works. In particular, for CelebA we use the accuracy of the underrepresented groups (i.e., bias-conflicting accuracy) and the average accuracy across all groups (i.e., unbiased accuracy) [39, 16]. For Waterbirds, we use the Worst-Group (WG) accuracy and the average accuracy across all groups. In the case of UrbanCars, we calculate the weighted average accuracy across groups, referred to as In-Distribution Accuracy (I.D. Acc), with weights determined by group representation ratios. Furthermore, I.D. Acc serves as a baseline for assessing accuracy drops related to background (BG Gap), co-occurring objects (CoObj Gap), and both background and co-occurring objects combined (BG+CoObj Gap). Lastly, for ImageNet9, we use 7 test set variations for evaluating the impact of background on model’s predictions: ORIGINAL, ONLY-BG-B (black object’s bounding box), ONLY-BG-T (impainted object’s bounding box), NO-FG Black (removed segmented object), ONLY-FG (black background), MIXED-RAND (random background of a random class), and MIXED-NEXT (random background of the next class).
4.2 Comparative analysis
In this subsection, we compare the performance of MAVias with other competitive BLU methods. Comprehensive results, including comparisons with BLA methods, are provided in the Appendix. Table 2 presents the performance of MAVias on CelebA with gender as target and the BlondHair as the biased attribute. The performance on unbiased and bias-conflicting samples indicates the ability of MAVias to reduce reliance on spurious correlations between the target and the corresponding biased attributes. Specifically, MAVias achieves 87.1% accuracy on bias-conflicting samples, surpassing state-of-the-art BLU methods such as FLAC-B [39] and LfF [31], while maintaining high unbiased accuracy (89.7%).
Table 2: Evaluation on CelebA with BlondHair and gender as biased attribute and target, respectively.
Methods Unbiased Bias-conflict
LfF [31] 84.2$\pm$0.3 81.2$\pm$1.4
SoftCon [16] 84.1 74.4
FLAC-B [39] 87.0$\pm$0.6 84.9$\pm$2.2
MAVias 89.7$\pm$0.6 87.1$\pm$1.7
Table 3: Evaluation on Waterbirds.
Method WG Acc. Avg. Acc.
JTT [26] 86.7$\pm$1.5 93.3$\pm$0.3
DISC [46] 88.7$\pm$0.4 93.8$\pm$0.7
MAVias 93.7$\pm$0.4 94.5$\pm$0.4
Table 3 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment. Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BLU methods (+5%), while slightly improving the average accuracy (+0.7%). The results on UrbanCars in Tab. 4 further validate the efficacy of MAVias in mitigating multiple biases introduced by co-occurring objects and background features. While the in-distribution accuracy (I.D. Acc) of MAVias is marginally lower than that of LfF [31] or Debian [23], it achieves substantial improvements in reducing the BG Gap (-4.1) and CoObj Gap (-2.4), as well as their combined effects (BG+CoObj Gap: -6.7).
Table 4: Evaluation on UrbanCars.
Method I.D. Acc BG Gap CoObj Gap BG+CoObj Gap
LfF [31] 97.2 -11.6 -18.4 -63.2
JTT [26] 95.9 -8.1 -13.3 -40.1
Debian [23] 98.0 -14.9 -10.5 -69.0
MAVias 92.8$\pm$0.8 -4.1$\pm$0.6 -2.4$\pm$1.4 -6.7$\pm$1.4
Table 5: Accuracy on ImageNet9 test sets. The reported results for the compared methods are based on reimplementations using code provided by the authors of [24].
Method MIXED-NEXT ($\uparrow$) MIXED-RAND ($\uparrow$) NO-FG ($\downarrow$) ONLY-BG-B ($\downarrow$) ONLY-BG-T ($\downarrow$) ONLY-FG ($\uparrow$) ORIGINAL ($\uparrow$)
Vanilla 82.66 $\pm$0.1 85.06 $\pm$0.0 64.16 $\pm$0.1 35.18 $\pm$0.1 44.74 $\pm$0.1 93.12 $\pm$0.0 97.69$\pm$0.0
LfF [31] 78.70 $\pm$0.1 81.47 $\pm$0.2 61.07 $\pm$0.1 34.82 $\pm$0.2 44.46 $\pm$0.0 88.99 $\pm$0.2 94.34 $\pm$0.2
JTT [26] 84.43 $\pm$0.1 86.16 $\pm$0.5 61.09 $\pm$2.0 32.04 $\pm$1.0 36.62 $\pm$4.7 92.09 $\pm$0.5 97.71 $\pm$0.1
Debian [23] 83.02 $\pm$0.4 85.64 $\pm$0.3 64.53$\pm$0.4 34.45$\pm$0.1 45.00 $\pm$0.6 93.06 $\pm$0.1 97.89 $\pm$0.1
MAVias 88.26$\pm$0.1 (+5.24) 89.64$\pm$0.2 (+3.48) 53.02$\pm$0.7 (+8.05) 21.83$\pm$0.4 (-10.21) 32.48$\pm$0.6 (-4.14) 91.90$\pm$0.4 (-1.22) 96.92$\pm$0.2 (-0.97)
(a) Vanilla
(b) MAVias
Figure 3: Logits for UrbanCars training samples belonging to groups defined by the urban car class and Background (BG) and Co-occurring Object (CoObj) biases.
In addition, we examine the logit space of the models to demonstrate the effect of MAVias on the behavior of the main model. Similar to the previous experiment, the urban car class of the UrbanCars dataset is used. As one may observe in Fig. 3 (a), the vanilla (biased) model increases the logit values proportionally with the bias, ranging from  4 to  7. This is an expected behavior, as the biases (i.e., spurious correlations) are easier to learn than the actual target and thus the model illustrates higher confidence (or logit values) for the bias-aligned samples. This phenomenon is referred to as Gradient Starvation [34] and can describe how bias is introduced to the models. Also, it is theoretically shown that methodologies penalizing high logit values can mitigate this phenomenon [34]. As shown in Fig. 3 (b), the interactions between the main model and the projection layer in the training procedure (defined by Eq. s (2) and (3)) successfully reduce the main model’s logits values proportionally to the bias in the data.
Apart from the typical training datasets designed for BLA methods, we also consider experiments on ImageNet9, a widely used CV dataset known for various tasks and equipped with multiple test sets specifically designed for bias evaluation. As shown in Tab. 5, MAVias achieves great improvements in the model’s accuracy on the test sets with modified backgrounds (i.e., MIXED-NEXT and MIXED-RAND), with the highest gains occurring for the MIXED-NEXT set (+5.24%) compared to the second best performing method. Similarly, for the sets where background information is suppressed (ONLY-BG-B, ONLY-BG-T, and NO-FG) MAVias achieves improvements of +10.21%, +4.14%, and +8.05%, respectively. As observed, all competitive methods, including MAVias, demonstrate reduced performance on ONLY-FG compared to the vanilla model. This is likely because the vanilla model can exploit biases present in the foreground (e.g., colors) to increase its accuracy. Finally, there is a 0.97% drop in performance on the ORIGINAL test set, which aligns with expectations, as MAVias is designed to rely less on shortcuts that boost overall performance. Furthermore, we evaluate MAVias using two subsets of ImageNet9 ORIGINAL, constructed based on tags identified by MAVias. Specifically, the bias-aligned set includes samples with tags for which the vanilla model performs above its overall accuracy, suggesting high bias impact; the rest of the samples form the bias-conflicting set. Table 6 demonstrates that MAVias improves accuracy on bias-conflicting samples by reducing reliance on visual biases, with a slight decrease on bias-aligned samples for the same reason. Finally, to gain a clearer understanding, Fig. 4 visualizes the bias-aligned tags for the Dog class in the bias-aligned set.
Table 6: Evaluation on the ImageNet9 ORIGINAL set comparing accuracy on the samples with irrelevant tags for which the Vanilla model overperforms (bias-aligned, BA) and the rest of the samples (bias-conflicting, BC).
Method BA acc. BC acc. Gap
Vanilla 98.55 $\pm$0.04 90.86 $\pm$0.11 -7.69 $\pm$0.16
MAVias 97.3 $\pm$0.21 94.13 $\pm$0.01 -3.17 $\pm$0.21
Figure 4: Visualization of the biases identified using MAVias for the Dog class in the ImageNet9 dataset.
4.3 Ablation Analysis
An important aspect in MAVias is how the embeddings are calculated. Table 7 compares two well-known vison-language models, i.e., CLIP and SigLip, and the text-based model BERT.
Table 7: Impact of the employed encoders on the MAVias performance. Results pertain to the Waterbirds dataset.
Model WG Acc. Avg. Acc.
BERT [7] 91.7$\pm$1.3 92.8$\pm$1.0
SigLip [48] 92.7$\pm$1.2 93.5$\pm$1.0
CLIP [36] 93.7$\pm$0.4 94.5$\pm$0.4
Furthermore, Tab. 8 compares MAVias’s performance using various LLMs. ChatGPT-4o shows the highest performance, with open-source models also performing competitively. For experiments, we use ChatGPT-4o to minimize LLM-related errors.
Table 8: Impact of the employed LLM on the MAVias performance. Results pertain to the Waterbirds dataset.
Model #parameters WG Acc. Avg. Acc.
Qwen2.5 [17] 7b 91.9$\pm$2.4 94.4$\pm$ 0.5
Mistral-small [18] 22b 92.9$\pm$0.4 94.9$\pm$0.2
Gemma2 [43] 9b 93.5$\pm$0.6 94.5$\pm$0.6
Llama3.1 [10] 8b 93.6$\pm$0.6 94.4$\pm$0.4
ChatGPT-4o [33] >>much-greater-than$>$ >175b 93.7$\pm$0.4 94.5$\pm$0.4
Finally, as discussed in Sec. 3, the regularization term in Eq.  (3) is crucial for balancing the logits between the two models. Table 9 shows the impact of removing this term from the training on the effectiveness of MAVias.
Table 9: Impact of Eq.  (3) on the MAVias performance. Results pertain to Waterbirds dataset.
Type WG Acc. Avg. Acc.
w/o regularizer 89.4$\pm$1.6 95.2$\pm$0.4
w/ regularizer 93.7$\pm$0.4 94.5$\pm$0.4
Note that in the Appendix, we also investigate different methods for computing embeddings from tags and examine the effect of the tags’ vocabulary size.
5 Conclusion
In this work, we presented a framework to address open-set biases in CV using instance-level descriptive tags. Unlike traditional methods relying on predefined biases, MAVias offers a flexible, scalable solution for identifying and mitigating unknown biases. It detects visual features that may introduce bias and reduces their impact through a robust approach. Extensive experiments show MAVias outperforms existing methods in detecting and mitigating complex, real-world biases. The main requirement is an image tagging model with a comprehensive tag vocabulary and an LLM suited to the data context. For example, cloud-deployed models are not ideal for sensitive data like medical images.
Appendix A Appendix
A.1 Theoretical Justification
A.1.1 Impact on Gradients
Let us consider a bias-aligned $\mathbf{x}^{(a)}$ and a bias-conflicting sample $\mathbf{x}^{(c)}$ with targets $y^{(a)}=y^{(c)}=\kappa$. The logits produced by $f_{\boldsymbol{\theta}}$ and $g_{\boldsymbol{\phi}}$ for each class $k\in\mathcal{Y}$, are then $\mathbf{z}^{(a)}_{\text{main}}(k)$, $\mathbf{z}^{(c)}_{\text{main}}(k)$, $\mathbf{z}^{(a)}_{\text{tag}}(k)$, and $\mathbf{z}^{(c)}_{\text{tag}}(k)$, respectively. $g_{\boldsymbol{\phi}}$ encodes visual bias, being irrelevant to class $\kappa$, thus we expect $\mathbf{z}^{(a)}_{\text{tag}}(\kappa)\gg\mathbf{z}^{(c)}_{\text{tag}}(\kappa)$. Also, bias-aligned samples that contain shortcuts are easier to learn by $f_{\boldsymbol{\theta}}$ and therefore it is expected that $\mathbf{z}^{(a)}_{\text{main}}(\kappa)\geq\mathbf{z}^{(c)}_{\text{main}}(\kappa)$. These imply:
$\begin{split}\mathbf{z}^{(a)}_{\text{main}}(\kappa)+\mathbf{z}^{(a)}_{\text{tag}}(\kappa)&\gg\mathbf{z}^{(c)}_{\text{main}}(\kappa)+\mathbf{z}^{(c)}_{\text{tag}}(\kappa)\Rightarrow\\ \sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})&\gg\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})\Rightarrow\\ \frac{1}{\sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})}&\ll\frac{1}{\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})}\end{split}$ (4)
where $\sigma^{(k)}$ denotes the $k$th class’ softmax score.
Additionally, $f_{\boldsymbol{\theta}}$ after a few iterations learns bias-aligned samples (based either on relevant or irrelevant class features) indicating $\operatorname*{arg\,max}_{k\in\mathcal{Y}}\mathbf{z}^{(a)}_{\text{main}}(k)=\kappa$, while $\operatorname*{arg\,max}_{k\in\mathcal{Y}}\mathbf{z}^{(a)}_{\text{tag}}(k)=\kappa$, by definition. Hence, $\sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})\approx 1$ leading to diminished gradients $\frac{\partial}{\partial\theta_{0}}[\sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})]\approx 0$. On the other hand, $\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})\in(0,1)$ leading to gradients $\frac{\partial}{\partial\theta_{0}}[\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})]>0$. Hence:
$\frac{\partial}{\partial\theta_{0}}[\sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})]\leq\frac{\partial}{\partial\theta_{0}}[\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})]$ (5)
Combining Eq. 4 and Eq. 5 we get:
$\begin{split}\frac{1}{\sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})}\cdot\frac{\partial}{\partial\theta_{0}}[\sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})]&\ll\\ \frac{1}{\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})}\cdot\frac{\partial}{\partial\theta_{0}}[\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})]&\Rightarrow\\ \frac{\partial}{\partial\theta_{0}}[\log{\sigma^{(\kappa)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})}]&\ll\\ \frac{\partial}{\partial\theta_{0}}[\log{\sigma^{(\kappa)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})}]&\Rightarrow\\ \frac{\partial}{\partial\theta_{0}}\bigg{[}\sum_{k}\mathbbm{1}(a,k)\cdot\log{\sigma^{(k)}(\mathbf{z}^{(a)}_{\text{main}}+\mathbf{z}^{(a)}_{\text{tag}})}\bigg{]}&\ll\\ \frac{\partial}{\partial\theta_{0}}\bigg{[}\sum_{k}\mathbbm{1}(c,k)\cdot\log{\sigma^{(k)}(\mathbf{z}^{(c)}_{\text{main}}+\mathbf{z}^{(c)}_{\text{tag}})}\bigg{]}&\Rightarrow\\ \frac{\partial\mathcal{L}^{cls}(\mathbf{x}^{(a)})}{\partial\theta_{0}}\ll\frac{\partial\mathcal{L}^{cls}(\mathbf{x}^{(c)})}{\partial\theta_{0}}\end{split}$ (6)
where
$\mathbbm{1}(i,k)=\begin{cases}1&\text{if }y_{i}=k\\ 0&\text{otherwise}\end{cases}$ (7)
which indicates that through our framework the main model $f_{\boldsymbol{\theta}}$ is enforced to update its weights based mainly on the gradients corresponding to the bias-conflicting samples $\mathbf{x}^{(c)}$. This essentially means that $f_{\boldsymbol{\theta}}$ will eventually ignore the visual cues prevalent in $\mathbf{x}^{(a)}$ samples corresponding to irrelevant information encoded by the tag model $g_{\boldsymbol{\phi}}$.
A.1.2 Unbiased Distribution Learning
Let $p_{tr}(X,Y,B)$ be the training data distribution where $X$, $Y$, $B$ are random variables associated with the data, the label and the bias attribute, respectively. From Theorem 1 of [16] we know that if $P_{u}$ is a shift of $p_{tr}$ such that the distribution $P_{u}(Y|B)$ is uniform (and therefore unbiased), then for a model $f_{\boldsymbol{\theta}}=p_{u}(y|\mathbf{x})$ that estimates the conditional probability over $P_{u}$, its probability over $p_{tr}$ will be
$p_{tr}(y|\mathbf{x})=\frac{\exp\left(\mathbf{z}_{\text{main}}(y)+\log p_{tr}(y|b)\right)}{\sum_{y^{\prime}}\exp\left(\mathbf{z}_{\text{main}}(y^{\prime})+\log p_{tr}(y^{\prime}|b)\right)}$ (8)
In the case of MAVias, let the probability $p_{tr}(y|b)$ be estimated through the projection and classification layers, $\mathbf{z}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(g_{\boldsymbol{\phi}}(\mathbf{e}))$ which use the visual bias embeddings $\mathbf{e}$ (encoding the irrelevant tags) to produce logits $\mathbf{z}_{\text{tag}}$. As a result,
$\begin{split}\log p_{tr}(y|b)&=\log\frac{e^{\mathbf{z}_{\text{tag}}(y)}}{\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}}\\ &=\mathbf{z}_{\text{tag}}(y)-\log\left({\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}}\right)\\ &=\mathbf{z}_{\text{tag}}(y)-A(\mathbf{z}_{\text{tag}})\end{split}$ (9)
where $A(\mathbf{z}_{\text{tag}})=\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}$. Replacing Eq. (9) to Eq. (8) gives
$\begin{split}&p_{tr}(y|\mathbf{x})=\\ =&\frac{\exp\left(\mathbf{z}_{\text{main}}(y)+\mathbf{z}_{\text{tag}}(y)-A(\mathbf{z}_{\text{tag}})\right)}{\sum_{y^{\prime}}\exp\left(\mathbf{z}_{\text{main}}(y^{\prime})+\mathbf{z}_{\text{tag}}(y^{\prime})-A(\mathbf{z}_{\text{tag}})\right)}\\ =&\frac{\exp\left(-A(\mathbf{z}_{\text{tag}})\right)\exp\left(\mathbf{z}_{\text{main}}(y)+\mathbf{z}_{\text{tag}}(y)\right)}{\exp\left(-A(\mathbf{z}_{\text{tag}})\right)\sum_{y^{\prime}}\exp\left(\mathbf{z}_{\text{main}}(y^{\prime})+\mathbf{z}_{\text{tag}}(y^{\prime})\right)}\\ =&\frac{e^{(\mathbf{z}_{\text{main}}(y)+\mathbf{z}_{\text{tag}}(y))}}{\sum_{y^{\prime}}e^{(\mathbf{z}_{\text{main}}(y^{\prime})+\mathbf{z}_{\text{tag}}(y^{\prime})})}\end{split}$ (10)
thus leading to the logit addition of Eq. (2).
For this approach to work, however, one must ensure that models $g_{\boldsymbol{\phi}}$ and $f_{\boldsymbol{\theta}_{c}}$ are successfully trained to produce logits $\mathbf{z}_{\text{tag}}$ that predict $p_{tr}(y|b)$ and are not dominated by $\mathbf{z}_{\text{main}}$. This is achieved by the regularization term of Eq. (3) which controls the magnitude differences of $\mathbf{z}_{\text{main}}$ and $\mathbf{z}_{\text{tag}}$.
A.2 Experimental Setup
A.2.1 Model architectures
For comparability purposes, we employ the same network architectures as those used in other works [39, 16, 2, 40]. In particular, for CelebA, we adopt the ResNet-18 architecture [14], for Waterbirds, UrbanCars, and ImageNet9 datasets we use Resnet-50 networks. In all experiments, the projection layer is a dense layer that gets vision-language embeddings as input and its output size is aligned to the feature size of the main model.
A.2.2 Implementation details
The SGD optimizer is employed for all datasets except for CelebA, where Adam optimizer is used. We use an initial learning rate of 0.001, which is divided by 10 every 1/3 of the training epochs. The weight decay is set to $10^{-4}$. The batch size is 128 for CelebA and 64 for Waterbirds, UrbanCars, and ImageNet9. Following previous works [16, 37, 24], we train the models for 40, 100, 300, and 40 for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively, For Waterbirds and UrbanCars, we do not use a learning rate scheduler. The parameters $(\alpha,\lambda)$ are (0.01, 0.5), (0.05, 0.6), (0.01, 0.4), and (0.001, 0.7) for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively. The optimal values of the hyperparameters are derived through a grid search. RAM [49], ChatGPT-4o [33], CLIP [36] are employed for image tagging, irrelevant tag filtering, and vision-language encoding, respectively. Experiments on CelebA, Waterbirds, and UrbanCars, were conducted on an NVIDIA A100 GPU, while experiments on Imagenet were conducted on a single NVIDIA RTX-4090 GPU. All experiments were repeated for 5 different random seeds.
A.2.3 Prompting
To ensure accurate tag classification, we provide the LLM with a detailed system prompt. Tags are processed in batches of 100, as testing has shown that longer lists can lead to some tags being overlooked. Since relevant tags are significantly fewer than irrelevant ones, we instruct the LLM to return only the relevant tags, allowing us to deduce the irrelevant ones. The exact system prompt is presented in Fig. 5
System Prompt I will provide you with the name of a target class and a large list of tags. Your task is to evaluate the tags and identify only those directly related to the target class. A tag is considered relevant if it describes or is an essential part of the object associated with the class name. This includes tags that refer to: physical components, defining features, inherent characteristics, and essential behaviors or functions of the object. For example, if the target class is “bee” tags like “insect”, “wing”, and “buzz” are relevant because they describe core aspects of what a bee is or does. Conversely, a tag is irrelevant if it refers to elements that are not an intrinsic part of the object. Irrelevant tags may include: background details, environmental context, colors (unless a defining characteristic), lighting, textures, other objects, or other non-essential contextual elements. For example, in the case of the class “bee”, tags like “sky”, “flower”, or “blue” would be irrelevant, as they describe the environment or background rather than the bee itself. Please output only the relevant tags in JSON format only (i.e., { relevant_tags: [the list of tags]}).
Figure 5: LLM system prompt for deriving the relevant tags.
A.3 Two-moon Distribution Experiment
To demonstrate the capability of MAVias in mitigating biased features, we extend the classic two-moon distribution into a 3-dimensional setting. In this scenario, we consider a dataset $\mathcal{D}=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{N}$, where each input $\mathbf{x}^{(i)}=(x^{(i)}_{1},x^{(i)}_{2},x^{(i)}_{3})$ consists of two relevant features $x^{(i)}_{1},x^{(i)}_{2}$ and an additional irrelevant feature $x^{(i)}_{3}$ that introduces bias. The target label $y^{(i)}$ should be determined based on the features $x^{(i)}_{1},x^{(i)}_{2}$, following the typical two-moon structure, while the feature $x^{(i)}_{3}$ represents the bias (e.g., samples are linearly separable with respect to $x^{(i)}_{3}$, even though it should not be used for classification).
In this setup, the main model $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})$ receives the full input $\mathbf{x}^{(i)}=(x^{(i)}_{1},x^{(i)}_{2},x^{(i)}_{3})$, while the projection layer $g_{\boldsymbol{\phi}}(x^{(i)}_{3})$ is provided only with the irrelevant feature $x^{(i)}_{3}$. The projection layer helps the system to explicitly account for this bias by incorporating $x^{(i)}_{3}$ in a controlled manner, while the main model is encouraged to focus on the relevant features $x^{(i)}_{1}$ and $x^{(i)}_{2}$ for classification. As shown in Fig. 6, MAVias enables the main model to effectively learn the underlying two-moon distribution based on the features $x^{(i)}_{1},x^{(i)}_{2}$, while the vanilla model relies only on the biased feature $x^{(i)}_{3}$, treating it as a shortcut that prevents the model from learning the proper distribution.
(a) Vanilla predictions on axes $x$ and $y$.
(b) Vanilla predictions on axes $x$ and $z$.
(c) MAVias predictions on axes $x$ and $y$.
(d) MAVias predictions on axes $x$ and $z$.
Figure 6: Two-moon problem on 3 dimensions. The distributions are linearly separable on axis $z$ (i.e., $z$ feature introduces bias), while the actual target is to learn the distributions defined by the features $x$ and $y$.
A.4 Comparative Analysis
This section provides a comprehensive comparative analysis involving both bias label-aware (BA) and bias label-unaware (BU) methodologies. As shown in Tab. 10, MAVias demonstrates competitive performance on CelebA compared to the best-performing BA method, i.e., -0.6% bias-conflict accuracy and -1.5% unbiased accuracy, even though MAVias has no access to bias-related information.
Table 10: Evaluation of the proposed method on CelebA with BlondHair and gender as biased attribute and target, respectively. Bold denotes the best-performing bias label-unaware (BU) method and underlined denotes the best-performing bias label-aware method.
Methods BU Bias
BlondHair
Unbiased Bias-conflict
LNL [19] 80.1$\pm$0.8 61.2$\pm$1.5
DI [45] 90.9$\pm$0.3 86.3$\pm$0.4
EnD [42] 86.9$\pm$1.0 76.4$\pm$1.9
BC-BB [16] 91.4$\pm$0.0 87.2$\pm$0.2
FairKL [2] 81.7$\pm$1.7 69.9$\pm$2.4
FLAC [39] 91.2$\pm$0.3 88.7$\pm$0.5
LfF [31] 84.2$\pm$0.3 81.2$\pm$1.4
SoftCon [16] 84.1 74.4
FLAC-B [39] 87.0$\pm$0.6 84.9$\pm$2.2
MAVias 89.7$\pm$0.6 87.1$\pm$1.7
Furthermore, Tab. 11 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment. Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BA methods, i.e., GroupDRO, BAdd, and DFR. Also, MAVias achieves this improvement while maintaining competitive average accuracy (+0.3%), which indicates that the reduction in bias did not come at the cost of overall performance. In contrast, BU methods like JTT and DISC, which can be directly compared with MAVias, show a larger gap in WG performance, indicating their limited effectiveness as a result of the lack of bias information (i.e., type of bias or bias labels), a limitation that MAVias successfully addresses.
Table 11: Evaluation on Waterbirds.
Method BU WG Acc. Avg. Acc.
GroupDro [37] 90.6$\pm$1.1 91.8$\pm$0.3
BAdd [40] 92.9$\pm$0.3 93.6$\pm$0.2
DFR [35] 92.9$\pm$0.2 94.2$\pm$0.4
JTT [26] 86.7$\pm$1.5 93.3$\pm$0.3
DISC [46] 88.7$\pm$0.4 93.8$\pm$0.7
MAVias 93.7$\pm$0.4 94.5$\pm$0.4
Finally, Tab. 12 presents the performance of the relevant methods on the UrbanCars dataset. This benchmark poses a greater challenge due to the presence of multiple biases, as reflected in the results, i.e., most BU methods perform significantly worse than BA methods. However, MAVias achieves competitive performance, with only 2%, 0.8%, and 2.8% drops in BG, CoObj, and BG+CoObj Gaps, respectively, compared to the best-performing BA methods.
Table 12: Evaluation on UrbanCars.
Method BU I.D. Acc BG Gap CoObj Gap BG+CoObj Gap
GroupDro [37] 91.6 -10.9 -3.6 -16.4
DFR [35] 89.7 -10.7 -6.9 -45.2
LLE [24] 96.7 -2.1 -2.7 -5.9
BAdd [40] 91.0$\pm$0.7 -4.3$\pm$0.4 -1.6$\pm$1.0 -3.9$\pm$0.4
LfF [31] 97.2 -11.6 -18.4 -63.2
JTT [26] 95.9 -8.1 -13.3 -40.1
Debian [23] 98.0 -14.9 -10.5 -69.0
MAVias 92.8$\pm$0.8 -4.1$\pm$0.6 -2.4$\pm$1.4 -6.7$\pm$1.4
A.5 Ablation Study
First, let us report the accuracy of $g_{\boldsymbol{\phi}}(\cdot)$ on the training samples of the urban car class from the UrbanCars dataset, which correlates the target classes with relevant backgrounds and co-occurring objects. As shown in Tab. 13, the $g_{\boldsymbol{\phi}}$ shows a clear distinction in its performance: its accuracy increases from 0% for bias-conflicting samples (i.e., those with a country background and co-occurring object) to 98.42% for bias-aligned samples (i.e., those with an urban background and co-occurring object). This result confirms that our tag-based approach produces embeddings that effectively capture bias, providing a foundation for the main model $f_{\boldsymbol{\theta}}(\cdot)$ to focus on learning unbiased features.
Table 13: MAVias: $g_{\boldsymbol{\theta}}(\cdot)$ accuracy for the UrbanCars training samples belonging to subgroups of urban car class with different biases.
BG CoObj #samples $g_{\boldsymbol{\theta}}(\cdot)$ acc.
Country Country 10 00.00
Country Urban 190 54.21
Urban Country 190 67.36
Urban Urban 3610 98.42
Furthermore, we investigate the impact of the vocabulary size employed for image tagging. The original vocabulary RAM uses has a size of 4585 words. Tab. 14 reports the performance of MAVias for different portions of the original vocabulary. Notably, using 30% or more of the original vocabulary yields highly effective models, which is due to the dataset’s bias (aquatic or terrestrial environments) that a variety of tags can capture.
Table 14: Performance of MAVias across varying portions of the original tag vocabulary. Results pertain to the Waterbirds dataset.
size WG Acc. Avg. Acc.
10% 78.1 87.1
20% 79.4 89.0
30% 91.1 92.0
40% 93.1 92.3
50% 93.8 93.2
60% 93.2 93.5
70% 92.9 94.2
80% 93.0 94.3
90% 93.6 94.5
100% 93.7 94.5
Moreover, Tab. 15 reports the performance of MAVias when computing individual embeddings for each tag and then averaging them compared to the one computing a single embedding for all tags combined. As one may observe, although both are highly effective, the latter can provide more representative features.
Table 15: Comparison of MAVias performance on Waterbirds using different approaches for computing embeddings from tags. “Separately” denotes computing individual embeddings for each tag and then averaging them, while “collectively” denotes computing a single embedding for all tags combined.
Type WG Acc. Avg. Acc.
separately 93.0$\pm$0.2 93.6$\pm$0.2
collectively 93.7$\pm$0.4 94.5$\pm$0.4
Finally, Fig. 7 illustrates the impact of different values of hyperparameters $\alpha$ and $\lambda$ on the MAVias performance.
Figure 7: Impact of hyperparameters $\alpha$ and $\lambda$ on the Worst-Group accuracy on the Waterbirds dataset.

### v2

@@
 Abstract
-Mitigating biases in computer vision models is an essential step towards the trustworthiness of artificial intelligence models.
-Existing bias mitigation methods focus on a small set of predefined biases, limiting their applicability in visual datasets where multiple, possibly unknown biases exist.
-To address this limitation, we introduce MAVias, an open-set bias mitigation approach leveraging foundation models to discover spurious associations between visual attributes and target classes.
-MAVias first captures a wide variety of visual features in natural language via a foundation image tagging model, and then leverages a large language model to select those visual features defining the target class, resulting in a set of language-coded potential visual biases.
-We then translate this set of potential biases into vision-language embeddings and introduce an in-processing bias mitigation approach to prevent the model from encoding information related to them.
-Our experiments on diverse datasets, including CelebA, Waterbirds, ImageNet, and UrbanCars, show that MAVias effectively detects and mitigates a wide range of biases in visual recognition tasks outperforming current state-of-the-art.
-Figure 1: MAVias identifies visual biases through foundational models that extract tags representing visual features and assess relevance to the target class.
+Mitigating biases in computer vision models is an essential step towards trustworthy artificial intelligence systems.
+Existing bias mitigation methods are limited to predefined biases, preventing their use in visual datasets where multiple, possibly unknown biases exist.
+To address this limitation, we introduce MAVias, an open-set bias mitigation approach that leverages foundation models to discover spurious associations between visual attributes and target classes.
+MAVias first captures a wide variety of visual features in natural language via a foundation image tagging model, and then leverages a large language model to select visual features that define the target class, resulting in a set of language-coded potential visual biases.
+It then translates these biases into vision-language embeddings and introduces an in-processing bias mitigation approach to prevent the model from encoding information related to them.
+Experiments on diverse datasets, including CelebA, Waterbirds, ImageNet, and UrbanCars, show that MAVias effectively detects and mitigates a wide range of biases in visual recognition tasks, outperforming current state-of-the-art.
+Figure 1: MAVias identifies instance-level potential visual biases through foundational models that extract tags representing visual features and assess relevance to the target class.
 Then, MAVias encodes these features within the vision-language space and integrates them into a bias-aware framework to train a model that is invariant to such visual biases.
 1 Introduction
-Computer Vision (CV) progress has been largely driven by Deep Learning (DL) advances [9, 15] and large-scale datasets [6, 25, 21], enabling models to learn complex patterns and visual features with impressive accuracy.
-However, alongside this progress, concerns have emerged about biases embedded in these models [32, 29, 11, 38, 41, 44] – often stemming from unintended correlations present in the training data [27, 13, 30, 22].
-The correlated attributes that are irrelevant, act as “shortcuts” and can significantly impact the model’s reliability and generalization [37, 12].
+Computer Vision (CV) progress has been largely driven by Deep Learning (DL) advances [dosovitskiy2020image, ho2020denoising] and large-scale datasets [deng2009imagenet, lin2014microsoft, krishna2017visual], enabling models to learn complex patterns and visual features with impressive accuracy.
+However, alongside this progress, concerns have emerged about biases embedded in these models [ntoutsi2020bias, mehrabi2021survey, fabbrizzi2022survey, sarridis2023towards, sarridis2024facex, wang2022revise] – often stemming from unintended correlations present in the training data [liu2024devil, gupta2024bias, melnychuk2024bounds, li2024be].
+The correlated attributes that are irrelevant act as “shortcuts” and can significantly impact the model’s reliability and generalization [ramos2025, sagawa2019distributionally, gupta2018robot].
 It is important to note that in the context of this paper, we define as visual bias any characteristic that does not contribute to defining the target class, which we refer to as “irrelevant”.
 To address this, several methodologies have been developed.
-Broadly, these fall into two categories: Bias Label-Aware (BLA) and Label-Unaware (BLU) methods.
-BLA methods leverage the annotations of the attributes introducing the biases to address them [19, 45, 42, 2].
-BLU methods focus on extracting bias pseudo-labels in cases of extreme biases where a bias-proxy model (or bias-capturing classifier) can be trained using the task’s target labels that closely align with the bias labels [16, 31, 39, 28].
-While effective in certain contexts, both BLA and BLU methods are limited in their applicability when there are multiple, complex, and possibly unknown biases. Common challenging scenarios include:
-Unknown biases.
-In large-scale, general-purpose CV datasets, such as ImageNet, biases may be difficult to identify and largely remain unknown, as they vary widely between different classes and are not prominent enough to allow training of a bias proxy model.
-For instance, in the ImageNet9 example presented in Tab.
+Broadly, these fall into two categories: Bias label-Aware (BA) and label-Unaware (BU) methods.
+BA methods leverage the annotations of attributes introducing the biases to address them [kim2019lnl, wang2020DI, tartaglione2021end, barbano2022fairkl].
+BU methods focus on extracting bias pseudo-labels in cases of extreme biases where a bias-proxy model (or bias-capturing classifier) can be trained using the task’s target labels that closely align with the bias labels [hong2021bb, nam2020LfF, sarridis2023flac, ludebiasing].
+While effective in certain contexts, both BA and BU methods have limited applicability when there are multiple, complex, and possibly unknown biases. Common challenging scenarios include:
+  • Unknown biases: In large general-purpose CV datasets, such as ImageNet, biases are difficult to identify and largely remain unknown, as they vary widely across different classes and are not prominent enough to allow training of bias proxy models.
+For instance, in the ImageNet9 example of Tab.
 1, a sample labeled as a dog could introduce biases related to the background scene (e.g., armchair, couch, pillow, red), the color of the dog (e.g., black and white), or accessories like a neckband.
-Potential biases beyond a predefined set. The CelebA example in Tab. 1, shows that beyond the hair color, additional biases may be present, such as clothing styles (e.g., business suit, tie).
-Poor representation of predefined bias.
-In some cases, biases are reduced to single labels, such as “rural background” in the UrbanCars dataset.
+  • Potential biases beyond a predefined set: The CelebA example in Tab. 1, shows that beyond the hair color, additional biases may be present, such as clothing styles (e.g., business suit, tie).
+  • Poor representation of predefined bias: In some cases, biases are reduced to single labels, such as “rural background” in the UrbanCars dataset.
 However, as shown in Tab.
-1, more nuanced descriptors at the instance level (e.g., path, tree, wood, forest, hydrant, red, etc.) provide a richer context for modeling bias.
-Existing BLA and BLU methods are not designed to mitigate such biases, leading to models that are biased and/or do not achieve their optimal generalization potential.
+1, more nuanced descriptors at the instance level (e.g., path, tree, wood, forest, hydrant, red, etc.) provide a richer representation of bias.
+Table 1: Examples of extracted tags for various datasets. Red color indicates the irrelevant tags ($\mathcal{B}^{(i)}$).
+Dataset ImageNet9 Waterbirds UrbanCars CelebA
+Target Class dog bird species car type hair color
+Bias Type unknown background background and objects gender
+Sample [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Extracted Tags armchair, black, chair, couch, dog, neckband, pillow, red, sit, white bamboo, bamboo forest, bird, blue, branch, green, hide, parrot, perch, sit, stand, stem, tree, yellow car, path, forest, hydrant, lush, park, red, road, sedan, silver, SUV, tree, white, wood black, business suit, dress, dress shirt, man, stand, stare, suit, tie, wear
+Existing BA and BU methods are not designed to mitigate such biases, leading to models that are biased and/or do not achieve their optimal generalization potential.
 To address this, Mitigate Any Visual bias (MAVias) introduces a flexible and scalable solution, capable of identifying and mitigating open-set biases in CV datasets.
 MAVias begins by extracting descriptive tags that capture various visual features, such as general-purpose objects, actions, scenes, and visual attributes.
-Note that a sufficiently large and comprehensive vocabulary can cover the vast majority of visual features present in CV datasets, enabling the method to be considered as operating in an open-set setting.
+Recent advancements in image tagging [zhang2024recognize] cover large and comprehensive vocabularies (i.e., >4,000>4,000 tags), effectively meeting the open-set requirements of MAVias.
 Then, these tags are processed by a Large Language Model (LLM) to identify which ones are irrelevant to each of the target classes, leading to a rich set of language-encoded visual biases, text descriptions of visual characteristics that are irrelevant to the classification task at hand.
-We then translate these biases into vision-language embeddings, project them to the main model’s backbone space and then to the classification layer.
-This projection layer is trained simultaneously alongside the main model, and during training, the output logits are a linear combination of the logits of the main model and those of the projection layer that capture visual biases.
+MAVias translates these biases into vision-language embeddings, projects them to the main model’s backbone space and then to the classification layer.
+This projection layer is trained simultaneously alongside the main model, and during training, the output logits are a linear combination of the logits of the main model and those of the projection layer that captures visual biases.
 This setup allows the main model to be exposed to the biased features – those representing irrelevant information to the target class – in a controlled way that leads to bias-invariant representations.
 Overall, MAVias provides an effective, end-to-end solution for identifying and mitigating biases in open-set scenarios.
-We evaluate the proposed method on several datasets involving single-attribute biases (CelebA, Waterbirds) as well as multi-attribute predefined biases (UrbanCars) demonstrating state-of-the-art performance.
+We evaluate the proposed method on several datasets involving single-attribute biases (CelebA, Waterbirds) as well as multi-attribute predefined biases (UrbanCars), demonstrating state-of-the-art performance.
 Furthermore, experiments were conducted on ImageNet9, involving unknown biases, where the suggested approach demonstrates significant gains (from 5.24% to 10.21%) in terms of accuracy compared to existing competitive approaches.
-The main contributions of this paper are the following: (i) a framework for identifying instance-specific open-set potential visual biases in CV datasets, (ii) a learning strategy that exploits foundation models to learn bias-invariant representations, and (iii) an extensive evaluation study including comparative and ablation experiments across 4 thematically diverse datasets demonstrating the effectiveness and general applicability of the proposed method, which outperforms the current state-of-the-art.
+The main contributions of this paper are the following:
+  • A framework for identifying instance-specific open-set potential visual biases in CV datasets.
+  • A learning strategy that exploits foundation models to learn bias-invariant representations that force the under-training model to avoid encoding any number of identified potential biases per training sample.
+  • An extensive evaluation study including 4 thematically diverse datasets demonstrating the effectiveness and general applicability of MAVias, which outperforms the state-of-the-art.
+MAVias implementation is available as part of the VB-Mitigator library [sarridis2025vbmitigator].
 2 Related Work
-Bias identification
-Several methods recently emerged that leverage text (such as captions, keywords, or tags) for bias detection, highlighting the potential of this approach in this domain.
-For instance, Say My Name (SaMyNa) [4] is an explainability method that tries to discover model biases through text-based pipelines.
-Similarly, the Bias-to-Text (B2T) [20] and Language-guided (Lg) [50] frameworks discover biases by extracting common keywords from the captions of misclassified images.
-These methods aim to identify the main source of bias (e.g., hair color in the CelebA example illustrated in Fig.
-1) rather than discovering any potentially biased visual characteristic in an open-set setting.
-Furthermore, OpenBias [8] introduces a framework for detecting biases in text-to-image generative models by leveraging LLMs to propose potential biases from captions.
-Although operating in a different domain (text-to-image generation), similarly to MAVias  OpenBias acknowledges the need for discovering potential biases in an open-set setting.
-However, it solely relies on an LLM to generate specific bias lists for a given caption, raising concerns about whether the LLM’s interpretation of biases aligns with human judgment.
-In contrast, MAVias takes a more structured and systematic approach; it evaluates each descriptive tag by querying the LLM to determine whether it is directly relevant to the target class using a detailed prompt (refer to the Appendix for system prompt details).
-Bias mitigation
-Recent efforts to mitigate bias in CV have led to the development of various methodologies, including methods with direct access to the labels of attributes introducing bias (i.e., BLA methods) [19, 45, 42, 16, 2, 37, 35, 24] and methods that do not take advantage of such labels but instead rely on deriving pseudo-labels (i.e., BLU methods) [3, 1, 31, 39, 46].
-For instance, Learning Not to Learn (LNL) [19] is a BLA method that discourages the model from predicting the attribute introducing bias, while Bias Contrastive-Bias Balance (BC-BB) [16] and FairKL [2] rely on the bias labels to enforce bias-neutral pairwise similarities between the samples using contrastive learning.
-Some methods have indirect access to the bias labels by utilizing bias-capturing classifiers trained on different datasets offering bias-related information explicitly [39, 40].
-Finally, other methods infer pseudo-labels from the biased vanilla model to identify the biases [31, 1, 5].
+Bias identification.
+Several recent methods leverage text (such as captions, keywords, or tags) for bias detection, highlighting the potential of this approach in this domain.
+For instance, Say My Name (SaMyNa) [ciranni2024say] is an explainability method that tries to discover model biases through text-based pipelines.
+Similarly, the Bias-to-Text (B2T) [kim2024keywords] and Language-guided (Lg) [zhao2024language] frameworks discover biases by extracting common keywords from the captions of misclassified images.
+These methods aim to identify the main source of bias rather than discovering potentially biased visual characteristics in an open-set setting.
+Furthermore, OpenBias [d2024openbias] is a framework for detecting biases in text-to-image generative models by leveraging LLMs to propose potential biases from captions.
+Although operating in a different domain (text-to-image generation), similarly to MAVias  OpenBias acknowledges the need for discovering biases in an open-set setting.
+However, it provides an LLM with text-only image descriptions asking for potential types of bias thus missing visual grounding of the depicted information, which could offer essential context and semantics.
+In contrast, MAVias takes a more structured and systematic approach; it evaluates each descriptive tag by querying the LLM to determine whether it is directly relevant to the target class using a detailed prompt (specified in the supp.
+material).
+Furthermore, bias identification methods typically use a vanilla model and validation data to infer specific biases, while MAVias focuses on defining irrelevant visual features a priori, and leveraging them in model training to mitigate them.
+While some open-set identification methods can provide bias labels to define subgroups for use by existing bias mitigation approaches [ciranni2024say], it is known that mitigation methods struggle with multi-attribute biases and cannot scale beyond single-attribute subgrouping [li2023whac].
+In contrast, MAVias uses instance-level irrelevant visual features without relying on dataset statistics (i.e., neither bias labels nor subgroups w.r.t.
+them are defined), allowing it to handle complex biases effectively.
+Bias mitigation.
+Recent bias mitigation methods include those with direct access to the labels of attributes introducing bias (i.e., BA methods) [kim2019lnl, wang2020DI, tartaglione2021end, hong2021bb, barbano2022fairkl, sagawa2019distributionally, qiu2023simple, li2023whac] and those that do not take advantage of such labels but instead rely on deriving pseudo-labels (i.e., BU methods) [cadene2019rubi, bahng2020rebias, nam2020LfF, sarridis2023flac, wu2023discover, yang2024identifying, venkataramani2024causal].
+For instance, Learning Not to Learn (LNL) [kim2019lnl] is a BA method that discourages the model from predicting the attribute introducing bias, while Bias Contrastive-Bias Balance (BC-BB) [hong2021bb] and FairKL [barbano2022fairkl] rely on the bias labels to enforce bias-neutral pairwise similarities between the samples using contrastive learning.
+Some methods have indirect access to the bias labels by utilizing bias-capturing classifiers trained on different datasets offering bias-related information explicitly [sarridis2023flac, sarridis2024badd].
+Finally, other methods infer pseudo-labels from the biased vanilla model to identify the biases [nam2020LfF, bahng2020rebias, clark2019LM].
 While these methods have been effective in mitigating biases, they primarily depend on predefined bias labels or pseudo-labels derived from biased models.
-On the contrary, this work investigates open-set scenarios and introduces a flexible bias mitigation methodology that can discover and handle multiple, diverse biases without requiring a unified bias-labeling system across all the dataset’s samples.
+On the contrary, this work explores open-set scenarios and introduces a flexible bias mitigation approach that can discover and handle multiple, diverse biases without requiring a bias-labeling system.
 3 Methodology
@@
 The goal is to train a DL model $f_{\boldsymbol{\theta}}:\mathcal{X}\to\mathcal{Y}$, parameterized by $\boldsymbol{\theta}$, to predict the target label $y^{(i)}$ given an image $\mathbf{x}^{(i)}$, while mitigating a set of potential biases, $\mathcal{B}^{(i)}$, present in $\mathbf{x}^{(i)}$ that may lead to biased predictions.
-In this context the term “potential biases” refers to all visual attributes present in $\mathcal{X}$ that are irrelevant to the target class, such as elements of the background of images.
-3.2 Methodology
-Table 1: Examples of extracted tags for various datasets. Red color indicates the irrelevant tags ($\mathcal{B}^{(i)}$).
-Dataset ImageNet9 Waterbirds UrbanCars CelebA
-Target Class dog bird species car type gender
-Bias Type unknown background background and objects hair colour
-Sample [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Extracted Tags armchair, black, chair, couch, dog, neckband, pillow, red, sit, white, bamboo, bamboo forest, bird, blue, branch, green, hide, parrot, perch, sit, stand, stem, tree, yellow, car, path, forest, hydrant, lush, park, red, road, sedan, silver, SUV, tree, white, wood, black, business suit, dress, dress shirt, man, stand, stare, suit, tie, wear,
+In this context, the term “potential biases” refers to all visual attributes present in $\mathcal{X}$ that are irrelevant to the target class, such as elements of the background of images or other so-called “shortcuts”.
+3.2 Method
 3.2.1 Language-driven Bias Modeling
-First, MAVias utilizes an image tagging model to extract descriptive tags that represent the visual information present in an image.
-In many general-purpose CV datasets, biases consist of visual information that can be described in text.
-Formally, for each image $\mathbf{x}^{(i)}$, we derive a set of tags $\mathcal{T}^{(i)}=\{t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{m}\}$, where $m$ represents the total number of tags.
-These capture various visual attributes, including colors, objects, backgrounds, and other features that can either describe the target class or introduce bias.
-The next step is to filter out tags that should not influence the decisions of a robust classifier.
+In many general-purpose CV datasets, biases manifest through visual information that can be described in text.
+MAVias, first, utilizes an image tagging model to extract tags that describe the visual information present in an image.
+Formally, for each image $\mathbf{x}^{(i)}$, we derive a set of tags $\mathcal{T}^{(i)}=\{t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{m_{i}}\}$, where $m_{i}$ represents the total number of tags for $i$-th sample.
+These capture various visual attributes, including colors, objects, backgrounds, and other features that can either describe the target class or potentially introduce bias.
+The next step is to filter out tags that should not influence the decisions of an unbiased classifier.
 To achieve this, we leverage an LLM to identify which tags are irrelevant to the target class $y^{(i)}$.
@@
 To ensure the LLM correctly identifies these irrelevant tags, we carefully design the prompt used for this task by providing precise instructions on what tag should be considered relevant (e.g., physical components, defining features, inherent characteristics, etc.) or not (e.g., background details, lighting, textures, other objects, etc.).
-Details on the prompt formulation process are provided in the Appendix.
+Details on the prompt formulation process are provided in the supp.
+material.
 Table 1 shows several examples of $\mathcal{B}^{(i)}$ for samples belonging to different datasets.
 For each set of irrelevant tags $\mathcal{B}^{(i)}$ associated with an image $\mathbf{x}^{(i)}$, we employ a vision-language model to generate a single embedding $\mathbf{e}^{(i)}\in\mathbb{R}^{d}$, where $d$ is the dimension of the embedding.
-This is produced using the prompt “a photo of $t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{k}$”, where $k\leq m$ is the number of irrelevant tags.
+This is produced using the prompt “a photo of $t^{(i)}_{1},t^{(i)}_{2},\dots,t^{(i)}_{k_{i}}$”, where $k_{i}\leq m_{i}$ is the number of irrelevant tags for the $i$-th sample.
 This aggregate embedding captures the combined information from all irrelevant tags, providing an instance-level representation of the biased features that could affect the model’s behavior.
-Comparison with alternative embedding approaches is provided in Appendix.
+Comparison with alternative embedding approaches is provided in Sec.
+4.6.
 Figure 2: Illustration of the proposed framework for mitigating any visual bias during model training.
@@
 3.2.2 Bias Mitigation
-We define the main model $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})$, a DL classifier composed of the following: (i) a backbone $f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})$, which extracts feature representations $\mathbf{h}^{(i)}\in\mathbb{R}^{r}$, where $r$ is the feature vector size; (ii) a classification head $f_{\boldsymbol{\theta}_{c}}(\mathbf{x}^{(i)})$ outputting the logits $\mathbf{z}^{(i)}_{\text{main}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{h}^{(i)})\in\mathbb{R}^{p}$, where $p$ is the number of classes.
-Thus, the overall main model is expressed as:
-$f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})=f_{\boldsymbol{\theta}_{c}}(f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})).$ (1)
+We define the main model $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})$, a DL classifier composed of the following: (i) a backbone $f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)})$, which extracts feature representations $\mathbf{h}^{(i)}\in\mathbb{R}^{r}$, where $r$ is the feature vector size; (ii) a classification head $f_{\boldsymbol{\theta}_{c}}(\mathbf{x}^{(i)})$ outputting the logits $\mathbf{z}^{(i)}_{\text{main}}=f_{\boldsymbol{\theta}_{c}}(\mathbf{h}^{(i)})\in\mathbb{R}^{p}$, where $p=\lvert\mathcal{Y}\rvert$.
+The overall main model is expressed as $f_{\boldsymbol{\theta}}(\mathbf{x}^{(i)})=f_{\boldsymbol{\theta}_{c}}(f_{\boldsymbol{\theta}_{bb}}(\mathbf{x}^{(i)}))$.
 In parallel, we introduce a projection layer $g_{\boldsymbol{\phi}}$, parameterized by $\boldsymbol{\phi}$, which takes the visual bias embeddings $\mathbf{e}^{(i)}$ as input and outputs embeddings $\mathbf{b}^{(i)}\in\mathbb{R}^{p}$.
@@
 $\mathbf{z}^{(i)}=\mathbf{z}^{(i)}_{\text{main}}+\mathbf{z}^{(i)}_{\text{tag}}$. (1)
-It is worth noting that bias often arises during DL model training because biased attributes in the training data are easier to learn and thus dominate gradient updates [34].
+It is worth noting that bias often arises during DL model training because biased attributes in the training data are easier to learn and thus dominate gradient updates [pezeshki2021gradient].
 To counteract this, MAVias incorporates visual bias logits into the main model’s logits, ensuring that as the bias in a sample increases, its impact on gradient updates is reduced.
@@
 This is supported both empirically in Sec.
-4.2 and theoretically in the Appendix.
+4.5 and theoretically in the supp.
+material.
 In other words, $\mathbf{z}_{tag}$ assists in decoupling the learning of biased features from the actual task at hand, which allows the main model to focus on the relevant features for the target prediction.
 Furthermore, since both the main model and the projection layer are trained concurrently, it is essential to ensure the stability of the training process.
-To achieve this, we introduce a loss function that combines the classification loss with a regularization term.
-The classification loss ensures that the combined logits $\mathbf{z}^{(i)}$ accurately predict the target label $y^{(i)}$, while the regularization term aligns the magnitude of the main model’s logits $\mathbf{z}^{(i)}_{\text{main}}$ with those of the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$.
+To achieve this, we introduce a loss function that combines the classification loss with a logit alignment term.
+The classification loss ensures that the combined logits $\mathbf{z}^{(i)}$ accurately predict the target label $y^{(i)}$, while the alignment term controls the relative magnitudes of the main model’s logits $\mathbf{z}^{(i)}_{\text{main}}$ and the visual bias logits $\mathbf{z}^{(i)}_{\text{tag}}$.
 By doing so, we prevent $g_{\boldsymbol{\phi}}$ from dominating or being overshadowed by $h_{\boldsymbol{\theta}}$.
-Specifically, the total loss is computed as:
-$\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})+\alpha\cdot\mathcal{L}_{reg}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$
-where: (i) $\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})$ is the classification loss (e.g., cross-entropy loss) between the final logits $\mathbf{z}^{(i)}$ and the ground truth label $y^{(i)}$; (ii) $\mathcal{L}_{reg}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$ is the regularization term for the norm of the logits, calculated as:
-$\mathcal{L}^{(i)}_{\text{reg}}=\frac{1}{2}\left\|\|\mathbf{z}^{(i)}_{\text{main}}\|-\lambda\cdot\|\mathbf{z}^{(i)}_{\text{tag}}\|\right\|^{2}$ (3)
+Specifically, the loss is computed as: $\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\phi})=\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})+\alpha\cdot\mathcal{L}_{align}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$, where: (i) $\mathcal{L}_{cls}(\mathbf{z}^{(i)},y^{(i)})$ is the classification loss (e.g., cross-entropy loss) between the final logits $\mathbf{z}^{(i)}$ and the ground truth label $y^{(i)}$; (ii) $\mathcal{L}_{align}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$ is the logit alignment term for the norm of the logits, calculated as:
+$\mathcal{L}^{(i)}_{align}=\frac{1}{2}\left\|\|\mathbf{z}^{(i)}_{\text{main}}\|-\lambda\cdot\|\mathbf{z}^{(i)}_{\text{tag}}\|\right\|^{2}$ (2)
+where $\|\cdot\|$ denotes the $l^{2}$-norm, $\lambda\in(0,1)$ is a scaling factor and $\alpha\in(0,1)$ is a weighting factor that balances the influence of the $\mathcal{L}_{align}(\mathbf{z}^{(i)}_{\text{main}},\mathbf{z}^{(i)}_{\text{tag}})$ in the total loss function.
+Typically, the greater the bias in the data, the smaller the value of $\lambda$ should be, as this leads to smaller $\mathbf{z}_{main}$ values and reduced gradient updates for the bias-aligned samples.
+Details on the effects of hyperparameters $\lambda$ and $\alpha$ are provided in the supp.
+material.
+The overview of the proposed framework is illustrated in Fig.
+2.
+4 Experiments
+4.1 Datasets
+We use Biased-CelebA [hong2021bb], Waterbirds [sagawa2019distributionally], UrbanCars [li2023whac], and Imagenet9 [xiao2021noise].
+Biased-CelebA is a subset of the CelebA dataset, containing facial images annotated with 40 binary attributes.
+In this subset, BlondHair is the target attribute, while the gender attribute introduces bias with a 90% correlation.
+Waterbirds features a 95% co-occurrence between waterbirds (or landbirds) and aquatic (or terrestrial) backgrounds.
+UrbanCars is an artificially generated dataset with a 95% co-occurrence between car body types and relevant urban or rural backgrounds.
+ImageNet9 consists of 9 coarse ImageNet classes.
+4.2 Model Architectures
+For comparability purposes, we employ the same network architectures as those used in other works [sarridis2023flac, hong2021bb, barbano2022fairkl, sarridis2024badd].
+In particular, for CelebA, we adopt the ResNet-18 architecture [he2016deep], for Waterbirds, UrbanCars, and ImageNet9 datasets, we use ResNet-50 networks.
+In all experiments, the projection layer is a dense layer that gets vision-language embeddings as input and its output size is aligned with the feature size of the main model.
+4.3 Implementation Details
+The SGD optimizer is employed for all datasets except for CelebA, where Adam optimizer is used.
+We use an initial learning rate of 0.001, which is divided by 10 every 1/3 of the training epochs.
+The weight decay is set to $10^{-4}$.
+The batch size is 128 for CelebA and 64 for Waterbirds, UrbanCars, and ImageNet9.
+Following previous works [hong2021bb, sagawa2019distributionally, li2023whac], we train the models for 40, 100, 300, and 40 for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively.
+For Waterbirds and UrbanCars, we do not use a learning rate scheduler.
+The parameters $(\alpha,\lambda)$ are (0.01, 0.5), (0.05, 0.6), (0.01, 0.4), and (0.001, 0.7) for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively.
+RAM [zhang2024recognize], GPT-4o [openai2023chatgpt4], OpenCLIP [clip] are employed for image tagging, irrelevant tag filtering, and vision-language encoding, respectively.
+For all the experiments presented in the main manuscript, we run the compared methods using the code provided by the corresponding papers, using the suggested hyperparameters.
+Specifically, we use 50 epochs for bias discovery and 100 upweight for JTT; $\alpha=110$ for FLAC-B; 0.1 SD coefficient; and $\alpha=0.7$ for LfF.
+For the baseline methods results presented in the supplementary material (i.e., closed-set), we present performance as reported in the corresponding works to avoid potential underperformance due to reimplementation.
+Experiments were conducted on an NVIDIA A100 GPU.
+All experiments were repeated for 5 different random seeds.
+Table 2: Open-set performance comparison across CelebA, Waterbirds, and UrbanCars datasets.
+Dataset Method WG Acc. Avg. Acc.
+CelebA
+LfF [nam2020LfF] 14.7 $\pm$15.2 67.1 $\pm$4.4
+JTT [liu2021just] 31.5 $\pm$8.0 61.6 $\pm$8.5
+SD [pezeshki2021gradient] 13.3 $\pm$8.2 67.4 $\pm$1.9
+Debian [li2022discover] 12.0 $\pm$8.7 67.0 $\pm$2.6
+FLAC-B [sarridis2023flac] 12.0 $\pm$8.7 65.9 $\pm$2.4
+MAVias 66.7 $\pm$4.7 81.4 $\pm$1.8
+Waterbirds
+LfF [nam2020LfF] 30.0 $\pm$6.8 72.7 $\pm$1.4
+JTT [liu2021just] 64.7 $\pm$2.4 85.2 $\pm$4.6
+SD [pezeshki2021gradient] 35.0 $\pm$16.3 75.5 $\pm$4.1
+Debian [li2022discover] 37.5 $\pm$0.0 74.7 $\pm$0.5
+FLAC-B [sarridis2023flac] 37.5 $\pm$8.8 75.2 $\pm$2.7
+MAVias 75.4 $\pm$0.9 87.5 $\pm$1.2
+UrbanCars
+LfF [nam2020LfF] 34.6 $\pm$2.6 61.0 $\pm$1.4
+JTT [liu2021just] 69.0 $\pm$3.3 77.8 $\pm$0.3
+SD [pezeshki2021gradient] 40.4 $\pm$2.7 66.5 $\pm$1.1
+Debian [li2022discover] 33.2 $\pm$8.0 61.1 $\pm$2.1
+FLAC-B [sarridis2023flac] 28.5 $\pm$4.3 57.3 $\pm$1.7
+MAVias 84.4 $\pm$2.2 89.3 $\pm$1.3
+Table 3: Open-set performance on ImageNet9 in terms of accuracy across 7 test sets.
+Method MIXED-NEXT ($\uparrow$) MIXED-RAND ($\uparrow$) NO-FG ($\downarrow$) ONLY-BG-B ($\downarrow$) ONLY-BG-T ($\downarrow$) ONLY-FG ($\uparrow$) ORIGINAL ($\uparrow$)
 […]
-4 Experiments
-4.1 Datasets and Evaluation Protocol
-We employ Biased-CelebA [16], Waterbirds [37], UrbanCars [24], and Imagenet9 [47].
-Biased-CelebA is a subset of the well-known CelebA dataset consisting of facial images annotated with 40 binary attributes.
-In this subset, gender is treated as the target attribute, while the BlondHair attribute introduces bias by enforcing a 90% correlation between the two.
-Waterbirds is a dataset that features a co-occurrence of 95% between waterbirds (or landbirds) and aquatic environments (or terrestrial environments) as background.
-UrbanCars is an artificially generated dataset featuring a 95% co-occurrence between car body types and background and/or certain objects relevant to urban or rural regions.
-ImageNet9 consists of 9 coarse ImageNet classes (dog, bird, vehicle, reptile, carnivore, insect, instrument, primate, and fish).
-Regarding the evaluation, we follow the protocols suggested by the dataset providers and previous works.
-In particular, for CelebA we use the accuracy of the underrepresented groups (i.e., bias-conflicting accuracy) and the average accuracy across all groups (i.e., unbiased accuracy) [39, 16].
-For Waterbirds, we use the Worst-Group (WG) accuracy and the average accuracy across all groups.
+LfF [nam2020LfF] 78.70 $\pm$0.1 81.47 $\pm$0.2 61.07 $\pm$0.1 34.82 $\pm$0.2 44.46 $\pm$0.0 88.99 $\pm$0.2 94.34 $\pm$0.2
+JTT [liu2021just] 84.43 $\pm$0.1 86.16 $\pm$0.5 61.09 $\pm$2.0 32.04 $\pm$1.0 36.62 $\pm$4.7 92.09 $\pm$0.5 97.71 $\pm$0.1
+Debian [li2022discover] 83.02 $\pm$0.4 85.64 $\pm$0.3 64.53 $\pm$0.4 34.45 $\pm$0.1 45.00 $\pm$0.6 93.06 $\pm$0.1 97.89 $\pm$0.1
+SD [pezeshki2021gradient] 87.56 $\pm$0.57 88.92 $\pm$0.74 62.60 $\pm$1.05 31.42 $\pm$2.93 40.81 $\pm$3.00 93.71 $\pm$0.71 98.16 $\pm$0.06
+FLAC-B [sarridis2023flac] 84.60 $\pm$0.46 86.62 $\pm$0.45 59.84 $\pm$1.67 29.71 $\pm$0.53 40.38 $\pm$1.28 92.72 $\pm$0.73 97.89 $\pm$0.09
+[…]
+4.4 Evaluation Protocol
+To assess bias, we primarily use worst-group accuracy (WG Acc.), which measures the accuracy of the least-performing group within a dataset, and average accuracy (Avg.
+Acc.), which is the mean accuracy across all groups and reflects overall model performance.
+The formation of these groups determines whether the evaluation supports a closed- or open-set bias scenario.
+Accordingly, we implement two distinct evaluation protocols.
+The first one is designed to align with our approach, focusing on open-set biases, while the second protocol adheres to established evaluation standards in the literature (i.e., closed-set scenario).
+Furthermore, since MAVias is designed for BU scenarios, we evaluate its performance against the following widely-employed and competitive BU methods: LfF [nam2020LfF], JTT [liu2021just], SD [pezeshki2021gradient], Debian [li2022discover], and FLAC-B [sarridis2023flac].
+Open-set.
+Here, we form groups based on the presence/ absence of the detected open-set bias attributes (i.e., no predefined biases are considered).
+To achieve that, we use a vanilla model and the potential biases $\mathcal{B}$ extracted through MAVias to identify the subset of tags $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ that actually introduce bias to the model.
+We define bias as occurring when the model exhibits increased accuracy on images that include a particular tag, relative to its overall accuracy across the entire dataset.
+After deriving $\mathcal{B}^{\prime}$ for each class within the datasets, we categorize each group of images belonging to a class into two sub-groups based on whether the samples contain at least one of the biased tags.
+This results in $2\times p$ groups.
+Subsequently, we measure the WG Acc.
+and Avg.
+Acc.
+across all groups.
+To ensure a fair assessment, the optimal training epoch for each method is selected based on the overall accuracy, without considering any information related to the biases.
+An exception to this protocol is ImageNet9.
+In the original ImageNet9 test set, we observed that for several classes, the subgroups without any biases contain very few samples (fewer than 4), making reliable evaluation difficult.
+We therefore use the official seven test set variations, which allow for a more comprehensive evaluation of the model’s reliance on factors beyond the target object: ORIGINAL, ONLY-BG-B (bounding box of a black object), ONLY-BG-T (bounding box of an inpainted object), NO-FG Black (segmented object removed), ONLY-FG (black background), MIXED-RAND (random background of a random class), and MIXED-NEXT (random background of the next class).
+Closed-set.
+Here, we follow the protocols suggested by the dataset providers and previous works for predefined biases.
+In particular, for CelebA we use the accuracy of the underrepresented groups (i.e., bias-conflicting accuracy) and the average accuracy across all groups (i.e., unbiased accuracy) [sarridis2023flac, hong2021bb].
+For Waterbirds, we use the WG Acc.
+and the Avg.
+Acc.
+across all groups.
 In the case of UrbanCars, we calculate the weighted average accuracy across groups, referred to as In-Distribution Accuracy (I.D.
@@
 Acc serves as a baseline for assessing accuracy drops related to background (BG Gap), co-occurring objects (CoObj Gap), and both background and co-occurring objects combined (BG+CoObj Gap).
-Lastly, for ImageNet9, we use 7 test set variations for evaluating the impact of background on model’s predictions: ORIGINAL, ONLY-BG-B (black object’s bounding box), ONLY-BG-T (impainted object’s bounding box), NO-FG Black (removed segmented object), ONLY-FG (black background), MIXED-RAND (random background of a random class), and MIXED-NEXT (random background of the next class).
 4.5 Comparative Analysis
-In this subsection, we compare the performance of MAVias with other competitive BLU methods.
-Comprehensive results, including comparisons with BLA methods, are provided in the Appendix.
-Table 2 presents the performance of MAVias on CelebA with gender as target and the BlondHair as the biased attribute.
-The performance on unbiased and bias-conflicting samples indicates the ability of MAVias to reduce reliance on spurious correlations between the target and the corresponding biased attributes.
-Specifically, MAVias achieves 87.1% accuracy on bias-conflicting samples, surpassing state-of-the-art BLU methods such as FLAC-B [39] and LfF [31], while maintaining high unbiased accuracy (89.7%).
-Table 2: Evaluation on CelebA with BlondHair and gender as biased attribute and target, respectively.
-Methods Unbiased Bias-conflict
-LfF [31] 84.2$\pm$0.3 81.2$\pm$1.4
-SoftCon [16] 84.1 74.4
-FLAC-B [39] 87.0$\pm$0.6 84.9$\pm$2.2
-MAVias 89.7$\pm$0.6 87.1$\pm$1.7
-Table 3: Evaluation on Waterbirds.
-Method WG Acc. Avg. Acc.
-JTT [26] 86.7$\pm$1.5 93.3$\pm$0.3
-DISC [46] 88.7$\pm$0.4 93.8$\pm$0.7
-MAVias 93.7$\pm$0.4 94.5$\pm$0.4
-Table 3 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment.
-Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BLU methods (+5%), while slightly improving the average accuracy (+0.7%).
-The results on UrbanCars in Tab.
-4 further validate the efficacy of MAVias in mitigating multiple biases introduced by co-occurring objects and background features.
-While the in-distribution accuracy (I.D.
-Acc) of MAVias is marginally lower than that of LfF [31] or Debian [23], it achieves substantial improvements in reducing the BG Gap (-4.1) and CoObj Gap (-2.4), as well as their combined effects (BG+CoObj Gap: -6.7).
-Table 4: Evaluation on UrbanCars.
-Method I.D. Acc BG Gap CoObj Gap BG+CoObj Gap
-LfF [31] 97.2 -11.6 -18.4 -63.2
-JTT [26] 95.9 -8.1 -13.3 -40.1
-Debian [23] 98.0 -14.9 -10.5 -69.0
-MAVias 92.8$\pm$0.8 -4.1$\pm$0.6 -2.4$\pm$1.4 -6.7$\pm$1.4
-Table 5: Accuracy on ImageNet9 test sets. The reported results for the compared methods are based on reimplementations using code provided by the authors of [24].
-Method MIXED-NEXT ($\uparrow$) MIXED-RAND ($\uparrow$) NO-FG ($\downarrow$) ONLY-BG-B ($\downarrow$) ONLY-BG-T ($\downarrow$) ONLY-FG ($\uparrow$) ORIGINAL ($\uparrow$)
-Vanilla 82.66 $\pm$0.1 85.06 $\pm$0.0 64.16 $\pm$0.1 35.18 $\pm$0.1 44.74 $\pm$0.1 93.12 $\pm$0.0 97.69$\pm$0.0
-LfF [31] 78.70 $\pm$0.1 81.47 $\pm$0.2 61.07 $\pm$0.1 34.82 $\pm$0.2 44.46 $\pm$0.0 88.99 $\pm$0.2 94.34 $\pm$0.2
-JTT [26] 84.43 $\pm$0.1 86.16 $\pm$0.5 61.09 $\pm$2.0 32.04 $\pm$1.0 36.62 $\pm$4.7 92.09 $\pm$0.5 97.71 $\pm$0.1
-Debian [23] 83.02 $\pm$0.4 85.64 $\pm$0.3 64.53$\pm$0.4 34.45$\pm$0.1 45.00 $\pm$0.6 93.06 $\pm$0.1 97.89 $\pm$0.1
-MAVias 88.26$\pm$0.1 (+5.24) 89.64$\pm$0.2 (+3.48) 53.02$\pm$0.7 (+8.05) 21.83$\pm$0.4 (-10.21) 32.48$\pm$0.6 (-4.14) 91.90$\pm$0.4 (-1.22) 96.92$\pm$0.2 (-0.97)
+Table 2 presents the open-set performance comparison across the CelebA, Waterbirds, and UrbanCars datasets.
+For CelebA, MAVias significantly outperforms competing methods by achieving a 66.7% (+35.2%) WG accuracy and 81.4% (+14%) Avg.
+accuracy.
+As expected, most existing BU methods exhibit poor performance, as they struggle to handle scenarios with multiple concurrent biases [li2023whac].
+For the Waterbirds, MAVias reaches a WG accuracy of 75.4% (+10.7%) and an Avg.
+accuracy of 87.5% (+2.3%), surpassing competing approaches by a notable margin.
+Similarly, in the UrbanCars dataset, MAVias achieves a WG accuracy of 84.4% (+15.4%) and an Avg.
+accuracy of 89.3% (+11.5%).
+For ImageNet9, as shown in Tab.
+3, MAVias achieves accuracy improvements on the test sets with modified backgrounds, MIXED-NEXT and MIXED-RAND (+0.7% and +0.72).
+Similarly, for the sets where background information is suppressed (ONLY-BG-B, ONLY-BG-T, and NO-FG) MAVias achieves improvements of +7.88%, +4.14%, and +6.82%, respectively.
+As observed, most competitive methods, including MAVias, underperform on ONLY-FG compared to the vanilla model.
+This is likely because the vanilla model can exploit biases present in the foreground (e.g., colors) to increase its accuracy.
+Finally, there is a 1.24% drop in performance on the ORIGINAL test set, which aligns with expectations, as MAVias is designed to rely less on shortcuts that boost overall performance.
+The top-10 biased tags for each Imagenet9 class are reported in Tab.
+4.
+Moreover, Tab.
+5 reports the top-10 irrelevant tags derived using MAVias for CelebA, Waterbirds, and UrbanCars datasets.
+For CelebA, MAVias successfully detects the primary bias (i.e., gender) while also revealing other biases associated with facial expressions (smile), shot types (selfies), and clothing items (e.g., dresses, business suits, or ties).
+In the case of Waterbirds, we observe that, in addition to background elements, waterbirds are predominantly depicted in flight, whereas landbirds are often observed perched on tree branches.
+Finally, for UrbanCars, as anticipated, the top irrelevant tags correspond to objects that are common in urban/rural environments.
+Table 4: Top-10 biased tags for each class of ImageNet9.
+Class Top-10 Irrelevant Tags
+Bird perch, sit, branch, tree, tree branch, water, blue, twig, brown, sky
+Carnivore stand, grass, stone, lush, fur, walk, lay, enclosure, tree, red
+Dog brown, white, stand, black, lush, stare, green, carpet, blanket, bed
+Fish sea, swim, water, underwater, aquarium, tank, coral reef, catch, man, yellow
+Insect green, plant, break, floor, sit, stem, leaf, close-up, white, yellow
+Instrument play, woman, ceiling, table, pillar, cloth, hang, band, tool, hand
+Primate tree, sit, branch, tree branch, stare, black, enclosure, brown, stone, log
+Reptile floor, stone, green, branch, tree, grass, tree branch, water, lay, sit
+Vehicle road, track, red, building, curb, equipment, travel, load, train track, railroad
+Table 5: Top-10 biased tags for each class of CelebA, Waterbirds, and UrbanCars.
+$\mathcal{D}$ Class Top-10 Irrelevant Tags
+CelebA Blonde woman, dress, pose, smile, girl, beautiful, curl, selfie, actor, carpet
+Non-Blonde wear, man, shirt, selfie, tie, black, stand, stare, dress shirt, business suit
+Waterbirds Landbird stand, perch, sit, tree, yellow, branch, floor, brown, forest, green
+Waterbird water, stand, white, sea, fly, lake, sky, ledge, pond, stone
+UrbanCars Country Car road, rural, animal, stand, white, jump, cow, lift, bull, highway
+Urban Car park, house exterior, house, red, building, flip, home, lift, white, jump
+For the closed-set evaluation, as reported in Tab.
+6 MAVias achieves +2.7% (89.7%) Unbiased and +2.2 (87.1%) Bias-conflict accuracy compared to the second-best performing method, on CelebA.
+On Waterbirds, MAVias attains 93.7% WG Acc., outperforming all compared BU methods (+5%), while slightly improving the average accuracy (+0.7%).
+On UrbanCars, MAVias significantly enhances performance by reducing the BG, CoObj, and BG+CoObj gaps to 4.1, 2.4, and 6.7, respectively, representing absolute improvements of 4, 8.1, and 33.4.
+The full closed-set results are provided in the suppl.
+material.
+Table 6: Brief closed-set performance comparison between MAVias and competitive BU methods.
+Dataset Metric Best BU Method MAVias
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 (a) Vanilla
@@
 Figure 3: Logits for UrbanCars training samples belonging to groups defined by the urban car class and Background (BG) and Co-occurring Object (CoObj) biases.
-In addition, we examine the logit space of the models to demonstrate the effect of MAVias on the behavior of the main model.
-Similar to the previous experiment, the urban car class of the UrbanCars dataset is used.
-As one may observe in Fig.
-3 (a), the vanilla (biased) model increases the logit values proportionally with the bias, ranging from  4 to  7.
-This is an expected behavior, as the biases (i.e., spurious correlations) are easier to learn than the actual target and thus the model illustrates higher confidence (or logit values) for the bias-aligned samples.
-This phenomenon is referred to as Gradient Starvation [34] and can describe how bias is introduced to the models.
-Also, it is theoretically shown that methodologies penalizing high logit values can mitigate this phenomenon [34].
+In addition, by examining the logit spaces, as shown in Fig.
+3 (a), the vanilla model increases logit values proportionally with the bias, ranging from  4 to  7.
+This is an expected behavior, as the biases are easier to learn than the actual target and thus the model illustrates higher confidence for the bias-aligned samples.
+This phenomenon is referred to as Gradient Starvation [pezeshki2021gradient] and describes how bias is introduced.
+Also, it is theoretically shown that methodologies penalizing high logit values can mitigate this phenomenon [pezeshki2021gradient].
 As shown in Fig.
 3 (b), the interactions between the main model and the projection layer in the training procedure (defined by Eq.
-s (2) and (3)) successfully reduce the main model’s logits values proportionally to the bias in the data.
-Apart from the typical training datasets designed for BLA methods, we also consider experiments on ImageNet9, a widely used CV dataset known for various tasks and equipped with multiple test sets specifically designed for bias evaluation.
-As shown in Tab.
-5, MAVias achieves great improvements in the model’s accuracy on the test sets with modified backgrounds (i.e., MIXED-NEXT and MIXED-RAND), with the highest gains occurring for the MIXED-NEXT set (+5.24%) compared to the second best performing method.
-Similarly, for the sets where background information is suppressed (ONLY-BG-B, ONLY-BG-T, and NO-FG) MAVias achieves improvements of +10.21%, +4.14%, and +8.05%, respectively.
-As observed, all competitive methods, including MAVias, demonstrate reduced performance on ONLY-FG compared to the vanilla model.
-This is likely because the vanilla model can exploit biases present in the foreground (e.g., colors) to increase its accuracy.
-Finally, there is a 0.97% drop in performance on the ORIGINAL test set, which aligns with expectations, as MAVias is designed to rely less on shortcuts that boost overall performance.
-Furthermore, we evaluate MAVias using two subsets of ImageNet9 ORIGINAL, constructed based on tags identified by MAVias.
-Specifically, the bias-aligned set includes samples with tags for which the vanilla model performs above its overall accuracy, suggesting high bias impact; the rest of the samples form the bias-conflicting set.
-Table 6 demonstrates that MAVias improves accuracy on bias-conflicting samples by reducing reliance on visual biases, with a slight decrease on bias-aligned samples for the same reason.
-Finally, to gain a clearer understanding, Fig.
-4 visualizes the bias-aligned tags for the Dog class in the bias-aligned set.
-Table 6: Evaluation on the ImageNet9 ORIGINAL set comparing accuracy on the samples with irrelevant tags for which the Vanilla model overperforms (bias-aligned, BA) and the rest of the samples (bias-conflicting, BC).
-Method BA acc. BC acc. Gap
-Vanilla 98.55 $\pm$0.04 90.86 $\pm$0.11 -7.69 $\pm$0.16
-MAVias 97.3 $\pm$0.21 94.13 $\pm$0.01 -3.17 $\pm$0.21
-Figure 4: Visualization of the biases identified using MAVias for the Dog class in the ImageNet9 dataset.
+(1) and (2)) reduce the main model’s logit values proportionally to the data bias.
 4.6 Ablation Analysis
-An important aspect in MAVias is how the embeddings are calculated. Table 7 compares two well-known vison-language models, i.e., CLIP and SigLip, and the text-based model BERT.
+An important aspect in MAVias is how the embeddings are calculated. Table 7 compares two well-known vision-language models, i.e., OpenCLIP and SigLip, and the text-based model BERT.
 Table 7: Impact of the employed encoders on the MAVias performance. Results pertain to the Waterbirds dataset.
 Model WG Acc. Avg. Acc.
-BERT [7] 91.7$\pm$1.3 92.8$\pm$1.0
-SigLip [48] 92.7$\pm$1.2 93.5$\pm$1.0
-CLIP [36] 93.7$\pm$0.4 94.5$\pm$0.4
+[…]
+[…]
+[…]
 Furthermore, Tab.
 8 compares MAVias’s performance using various LLMs.
-ChatGPT-4o shows the highest performance, with open-source models also performing competitively.
-For experiments, we use ChatGPT-4o to minimize LLM-related errors.
+GPT-4o shows the highest performance, with open-source models also performing competitively.
+For experiments, we use GPT-4o to minimize LLM-related errors.
+Table 9 presents the performance of GPT-4o in deriving relevant tags for the Waterbirds, CelebA, and UrbanCars datasets.
+The ground truth was manually established.
 Table 8: Impact of the employed LLM on the MAVias performance. Results pertain to the Waterbirds dataset.
 Model #parameters WG Acc. Avg. Acc.
-Qwen2.5 [17] 7b 91.9$\pm$2.4 94.4$\pm$ 0.5
-Mistral-small [18] 22b 92.9$\pm$0.4 94.9$\pm$0.2
-Gemma2 [43] 9b 93.5$\pm$0.6 94.5$\pm$0.6
-Llama3.1 [10] 8b 93.6$\pm$0.6 94.4$\pm$0.4
-ChatGPT-4o [33] >>much-greater-than$>$ >175b 93.7$\pm$0.4 94.5$\pm$0.4
-Finally, as discussed in Sec.
-3, the regularization term in Eq.
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 9: Performance of GPT-4o.
+Dataset Precision Recall
+Waterbirds 96.1 79.0
+CelebA 81.8 75.0
+UrbanCars 89.2 71.7
+Furthermore, we investigate the impact of the different MAVias components.
+First, as discussed in Sec.
+3, the logit alignment term in Eq.
 (2) is crucial for balancing the logits between the two models.
@@
 Type WG Acc. Avg. Acc.
-w/o regularizer 89.4$\pm$1.6 95.2$\pm$0.4
-w/ regularizer 93.7$\pm$0.4 94.5$\pm$0.4
-Note that in the Appendix, we also investigate different methods for computing embeddings from tags and examine the effect of the tags’ vocabulary size.
+[…]
+[…]
+In addition, let us report the accuracy of $g_{\boldsymbol{\phi}}(\cdot)$ on the training samples of the urban car class from the UrbanCars dataset, which correlates the target classes with relevant backgrounds and co-occurring objects.
+As shown in Tab.
+11, the $g_{\boldsymbol{\phi}}$ shows a clear distinction in its performance: its accuracy increases from 0% for bias-conflicting samples (i.e., those with a country background and co-occurring object) to 98.42% for bias-aligned samples (i.e., those with an urban background and co-occurring object).
+This result confirms that our tag-based approach produces embeddings that effectively capture bias, providing a foundation for the main model $f_{\boldsymbol{\theta}}(\cdot)$ to focus on learning unbiased features.
+Table 11: MAVias: $g_{\boldsymbol{\theta}}(\cdot)$ accuracy for UrbanCars training samples belonging to subgroups of urban car class with different biases.
+BG CoObj #samples $g_{\boldsymbol{\theta}}(\cdot)$ acc.
+Country Country 10 00.00
+Country Urban 190 54.21
+Urban Country 190 67.36
+Urban Urban 3610 98.42
+Furthermore, we investigate the impact of the vocabulary size employed for image tagging.
+The original vocabulary of RAM has a size of 4585 words.
+Tab.
+12 reports the performance of MAVias for different portions of the original vocabulary.
+Notably, using 30% or more of the original vocabulary yields highly effective models.
+Table 12: Performance of MAVias across varying portions of the original tag vocabulary. Results pertain to the Waterbirds dataset.
+size WG Acc. Avg. Acc.
+10% 78.1 87.1
+20% 79.4 89.0
+30% 91.1 92.0
+40% 93.1 92.3
+50% 93.8 93.2
+60% 93.2 93.5
+70% 92.9 94.2
+80% 93.0 94.3
+90% 93.6 94.5
+100% 93.7 94.5
 5 Conclusion
-In this work, we presented a framework to address open-set biases in CV using instance-level descriptive tags.
-Unlike traditional methods relying on predefined biases, MAVias offers a flexible, scalable solution for identifying and mitigating unknown biases.
-It detects visual features that may introduce bias and reduces their impact through a robust approach.
+We presented a method to address open-set biases in CV using instance-level descriptive tags.
+Unlike previous methods relying on predefined biases, MAVias offers a flexible solution for identifying and mitigating unknown biases.
+It detects visual features that may introduce bias and reduces their impact.
 Extensive experiments show MAVias outperforms existing methods in detecting and mitigating complex, real-world biases.
-The main requirement is an image tagging model with a comprehensive tag vocabulary and an LLM suited to the data context.
-For example, cloud-deployed models are not ideal for sensitive data like medical images.
-Appendix A Appendix
-A.1 Theoretical Justification
-A.1.1 Impact on Gradients
+The main requirement is an image tagging model with a comprehensive tag vocabulary and an LLM suited to the application context.
+For example, cloud-deployed models are not ideal for applications where sensitive data is involved, e.g., medical images, images of personal identity documents, etc.
+Acknowledgments
+This research was supported by the EU Horizon Europe projects MAMMOth (grant no. 101070285), ELIAS (grant no. 101120237), and ELLIOT (grant no. 101214398).
+[…]
+Supplementary Material
+1 Theoretical Justification
+1.1 Impact on Gradients
 Let us consider a bias-aligned $\mathbf{x}^{(a)}$ and a bias-conflicting sample $\mathbf{x}^{(c)}$ with targets $y^{(a)}=y^{(c)}=\kappa$.
@@
 $\mathbbm{1}(i,k)=\begin{cases}1&\text{if }y_{i}=k\\ 0&\text{otherwise}\end{cases}$ (6)
-which indicates that through our framework the main model $f_{\boldsymbol{\theta}}$ is enforced to update its weights based mainly on the gradients corresponding to the bias-conflicting samples $\mathbf{x}^{(c)}$.
+which indicates that through our framework, the main model $f_{\boldsymbol{\theta}}$ is enforced to update its weights based mainly on the gradients corresponding to the bias-conflicting samples $\mathbf{x}^{(c)}$.
 This essentially means that $f_{\boldsymbol{\theta}}$ will eventually ignore the visual cues prevalent in $\mathbf{x}^{(a)}$ samples corresponding to irrelevant information encoded by the tag model $g_{\boldsymbol{\phi}}$.
-A.1.2 Unbiased Distribution Learning
-Let $p_{tr}(X,Y,B)$ be the training data distribution where $X$, $Y$, $B$ are random variables associated with the data, the label and the bias attribute, respectively.
-From Theorem 1 of [16] we know that if $P_{u}$ is a shift of $p_{tr}$ such that the distribution $P_{u}(Y|B)$ is uniform (and therefore unbiased), then for a model $f_{\boldsymbol{\theta}}=p_{u}(y|\mathbf{x})$ that estimates the conditional probability over $P_{u}$, its probability over $p_{tr}$ will be
+1.2 Unbiased Distribution Learning
+Let $p_{tr}(X,Y,B)$ be the training data distribution where $X$, $Y$, $B$ are random variables associated with the data, the label, and the bias attribute, respectively.
+From Theorem 1 of [hong2021bb] we know that if $P_{u}$ is a shift of $p_{tr}$ such that the distribution $P_{u}(Y|B)$ is uniform (and therefore unbiased), then for a model $f_{\boldsymbol{\theta}}=p_{u}(y|\mathbf{x})$ that estimates the conditional probability over $P_{u}$, its probability over $p_{tr}$ will be
 $p_{tr}(y|\mathbf{x})=\frac{\exp\left(\mathbf{z}_{\text{main}}(y)+\log p_{tr}(y|b)\right)}{\sum_{y^{\prime}}\exp\left(\mathbf{z}_{\text{main}}(y^{\prime})+\log p_{tr}(y^{\prime}|b)\right)}$ (7)
-In the case of MAVias, let the probability $p_{tr}(y|b)$ be estimated through the projection and classification layers, $\mathbf{z}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(g_{\boldsymbol{\phi}}(\mathbf{e}))$ which use the visual bias embeddings $\mathbf{e}$ (encoding the irrelevant tags) to produce logits $\mathbf{z}_{\text{tag}}$.
+In the case of MAVias, let the probability $p_{tr}(y|b)$ be estimated through the projection and classification layers, $\mathbf{z}_{\text{tag}}=f_{\boldsymbol{\theta}_{c}}(g_{\boldsymbol{\phi}}(\mathbf{e}))$, which use the visual bias embeddings $\mathbf{e}$ (encoding the irrelevant tags) to produce logits $\mathbf{z}_{\text{tag}}$.
 As a result,
 $\begin{split}\log p_{tr}(y|b)&=\log\frac{e^{\mathbf{z}_{\text{tag}}(y)}}{\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}}\\ &=\mathbf{z}_{\text{tag}}(y)-\log\left({\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}}\right)\\ &=\mathbf{z}_{\text{tag}}(y)-A(\mathbf{z}_{\text{tag}})\end{split}$ (8)
-where $A(\mathbf{z}_{\text{tag}})=\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}$. Replacing Eq. (9) to Eq. (8) gives
+where $A(\mathbf{z}_{\text{tag}})=\sum_{y^{\prime\prime}}e^{\mathbf{z}_{\text{tag}}(y^{\prime\prime})}$. Replacing (8) to (7) gives
 $\begin{split}&p_{tr}(y|\mathbf{x})=\\ =&\frac{\exp\left(\mathbf{z}_{\text{main}}(y)+\mathbf{z}_{\text{tag}}(y)-A(\mathbf{z}_{\text{tag}})\right)}{\sum_{y^{\prime}}\exp\left(\mathbf{z}_{\text{main}}(y^{\prime})+\mathbf{z}_{\text{tag}}(y^{\prime})-A(\mathbf{z}_{\text{tag}})\right)}\\ =&\frac{\exp\left(-A(\mathbf{z}_{\text{tag}})\right)\exp\left(\mathbf{z}_{\text{main}}(y)+\mathbf{z}_{\text{tag}}(y)\right)}{\exp\left(-A(\mathbf{z}_{\text{tag}})\right)\sum_{y^{\prime}}\exp\left(\mathbf{z}_{\text{main}}(y^{\prime})+\mathbf{z}_{\text{tag}}(y^{\prime})\right)}\\ =&\frac{e^{(\mathbf{z}_{\text{main}}(y)+\mathbf{z}_{\text{tag}}(y))}}{\sum_{y^{\prime}}e^{(\mathbf{z}_{\text{main}}(y^{\prime})+\mathbf{z}_{\text{tag}}(y^{\prime})})}\end{split}$ (9)
@@
 For this approach to work, however, one must ensure that models $g_{\boldsymbol{\phi}}$ and $f_{\boldsymbol{\theta}_{c}}$ are successfully trained to produce logits $\mathbf{z}_{\text{tag}}$ that predict $p_{tr}(y|b)$ and are not dominated by $\mathbf{z}_{\text{main}}$.
-This is achieved by the regularization term of Eq.
+This is achieved by the logit alignment term of Eq.
 (2) which controls the magnitude differences of $\mathbf{z}_{\text{main}}$ and $\mathbf{z}_{\text{tag}}$.
-A.2 Experimental Setup
-A.2.1 Model architectures
-For comparability purposes, we employ the same network architectures as those used in other works [39, 16, 2, 40].
-In particular, for CelebA, we adopt the ResNet-18 architecture [14], for Waterbirds, UrbanCars, and ImageNet9 datasets we use Resnet-50 networks.
-In all experiments, the projection layer is a dense layer that gets vision-language embeddings as input and its output size is aligned to the feature size of the main model.
-A.2.2 Implementation details
-The SGD optimizer is employed for all datasets except for CelebA, where Adam optimizer is used.
-We use an initial learning rate of 0.001, which is divided by 10 every 1/3 of the training epochs.
-The weight decay is set to $10^{-4}$.
-The batch size is 128 for CelebA and 64 for Waterbirds, UrbanCars, and ImageNet9.
-Following previous works [16, 37, 24], we train the models for 40, 100, 300, and 40 for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively, For Waterbirds and UrbanCars, we do not use a learning rate scheduler.
-The parameters $(\alpha,\lambda)$ are (0.01, 0.5), (0.05, 0.6), (0.01, 0.4), and (0.001, 0.7) for CelebA, Waterbirds, UrbanCars, and ImageNet9, respectively.
-The optimal values of the hyperparameters are derived through a grid search.
-RAM [49], ChatGPT-4o [33], CLIP [36] are employed for image tagging, irrelevant tag filtering, and vision-language encoding, respectively.
-Experiments on CelebA, Waterbirds, and UrbanCars, were conducted on an NVIDIA A100 GPU, while experiments on Imagenet were conducted on a single NVIDIA RTX-4090 GPU.
-All experiments were repeated for 5 different random seeds.
-A.2.3 Prompting
+2 Experimental Setup
+2.1 Prompting
 To ensure accurate tag classification, we provide the LLM with a detailed system prompt.
@@
 The exact system prompt is presented in Fig.
-5
+4
 System Prompt I will provide you with the name of a target class and a large list of tags.
@@
 This includes tags that refer to: physical components, defining features, inherent characteristics, and essential behaviors or functions of the object.
-For example, if the target class is “bee” tags like “insect”, “wing”, and “buzz” are relevant because they describe core aspects of what a bee is or does.
+For example, if the target class is “bee”, tags like “insect”, “wing”, and “buzz” are relevant because they describe core aspects of what a bee is or does.
 Conversely, a tag is irrelevant if it refers to elements that are not an intrinsic part of the object.
@@
 Figure 4: LLM system prompt for deriving the relevant tags.
-A.3 Two-moon Distribution Experiment
+3 Two-moon Distribution Experiment
 To demonstrate the capability of MAVias in mitigating biased features, we extend the classic two-moon distribution into a 3-dimensional setting.
@@
 As shown in Fig.
-6, MAVias enables the main model to effectively learn the underlying two-moon distribution based on the features $x^{(i)}_{1},x^{(i)}_{2}$, while the vanilla model relies only on the biased feature $x^{(i)}_{3}$, treating it as a shortcut that prevents the model from learning the proper distribution.
-(a) Vanilla predictions on axes $x$ and $y$.
-(b) Vanilla predictions on axes $x$ and $z$.
-(c) MAVias predictions on axes $x$ and $y$.
-(d) MAVias predictions on axes $x$ and $z$.
+5, MAVias enables the main model to effectively learn the underlying two-moon distribution based on the features $x^{(i)}_{1},x^{(i)}_{2}$, while the vanilla model relies only on the biased feature $x^{(i)}_{3}$, treating it as a shortcut that prevents the model from learning the proper distribution.
+(a) Vanilla outputs on axes $x$ and $y$.
+(b) Vanilla outputs on axes $x$ and $z$.
+(c) MAVias outputs on axes $x$ and $y$.
+(d) MAVias outputs on axes $x$ and $z$.
 Figure 5: Two-moon problem on 3 dimensions.
 The distributions are linearly separable on axis $z$ (i.e., $z$ feature introduces bias), while the actual target is to learn the distributions defined by the features $x$ and $y$.
-A.4 Comparative Analysis
-This section provides a comprehensive comparative analysis involving both bias label-aware (BA) and bias label-unaware (BU) methodologies.
-As shown in Tab.
-10, MAVias demonstrates competitive performance on CelebA compared to the best-performing BA method, i.e., -0.6% bias-conflict accuracy and -1.5% unbiased accuracy, even though MAVias has no access to bias-related information.
-Table 10: Evaluation of the proposed method on CelebA with BlondHair and gender as biased attribute and target, respectively.
+4 Comparative Analysis
+In this section, we compare the performance of MAVias with other competitive methods on closed-set (i.e., predefined bias) scenarios.
+The reported results encompass both bias-label aware (BA) and bias-label unaware (BU) methods.
+However, it should be noted that MAVias can be directly compared only with the BU approaches.
+Table 13 presents the performance of MAVias on CelebA with gender as target and the BlondHair as the biased attribute.
+The performance on unbiased and bias-conflicting samples indicates the ability of MAVias to reduce reliance on spurious correlations between the target and the corresponding biased attributes.
+Specifically, MAVias achieves 87.1% accuracy on bias-conflicting samples, surpassing state-of-the-art BU methods such as FLAC-B [sarridis2023flac] and LfF [nam2020LfF], while maintaining high unbiased accuracy (89.7%).
+Table 14 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment.
+Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BU methods (+5%), while slightly improving the average accuracy (+0.7%).
+The results on UrbanCars in Tab.
+15 further validate the efficacy of MAVias in mitigating multiple biases introduced by co-occurring objects and background features.
+While the in-distribution accuracy (I.D.
+Acc) of MAVias is marginally lower than that of LfF [nam2020LfF] or Debian [li2022discover], it achieves substantial improvements in reducing the BG Gap (4.1) and CoObj Gap (2.4), as well as their combined effects (BG+CoObj Gap: 6.7).
+Table 13: Bias-conflict and Unbiased accuracy comparison on CelebA with BlondHair and gender as biased attribute and target, respectively.
 Bold denotes the best-performing bias label-unaware (BU) method and underlined denotes the best-performing bias label-aware method.
@@
 Unbiased Bias-conflict
-LNL [19] 80.1$\pm$0.8 61.2$\pm$1.5
-DI [45] 90.9$\pm$0.3 86.3$\pm$0.4
-EnD [42] 86.9$\pm$1.0 76.4$\pm$1.9
-BC-BB [16] 91.4$\pm$0.0 87.2$\pm$0.2
-FairKL [2] 81.7$\pm$1.7 69.9$\pm$2.4
-FLAC [39] 91.2$\pm$0.3 88.7$\pm$0.5
-LfF [31] 84.2$\pm$0.3 81.2$\pm$1.4
-SoftCon [16] 84.1 74.4
-FLAC-B [39] 87.0$\pm$0.6 84.9$\pm$2.2
-MAVias 89.7$\pm$0.6 87.1$\pm$1.7
-Furthermore, Tab.
-11 evaluates MAVias on Waterbirds, a well-known benchmark for spurious correlation problems between bird type and background environment.
-Notably, MAVias attains 93.7% Worst-Group accuracy, outperforming all compared BA methods, i.e., GroupDRO, BAdd, and DFR.
-Also, MAVias achieves this improvement while maintaining competitive average accuracy (+0.3%), which indicates that the reduction in bias did not come at the cost of overall performance.
-In contrast, BU methods like JTT and DISC, which can be directly compared with MAVias, show a larger gap in WG performance, indicating their limited effectiveness as a result of the lack of bias information (i.e., type of bias or bias labels), a limitation that MAVias successfully addresses.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+SoftCon [hong2021bb] 84.1 74.4
+[…]
+[…]
 Table 14: Evaluation on Waterbirds.
 Method BU WG Acc. Avg. Acc.
-GroupDro [37] 90.6$\pm$1.1 91.8$\pm$0.3
-BAdd [40] 92.9$\pm$0.3 93.6$\pm$0.2
-DFR [35] 92.9$\pm$0.2 94.2$\pm$0.4
-JTT [26] 86.7$\pm$1.5 93.3$\pm$0.3
-DISC [46] 88.7$\pm$0.4 93.8$\pm$0.7
-MAVias 93.7$\pm$0.4 94.5$\pm$0.4
-Finally, Tab.
-12 presents the performance of the relevant methods on the UrbanCars dataset.
-This benchmark poses a greater challenge due to the presence of multiple biases, as reflected in the results, i.e., most BU methods perform significantly worse than BA methods.
-However, MAVias achieves competitive performance, with only 2%, 0.8%, and 2.8% drops in BG, CoObj, and BG+CoObj Gaps, respectively, compared to the best-performing BA methods.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 Table 15: Evaluation on UrbanCars.
 Method BU I.D. Acc BG Gap CoObj Gap BG+CoObj Gap
-GroupDro [37] 91.6 -10.9 -3.6 -16.4
-DFR [35] 89.7 -10.7 -6.9 -45.2
-LLE [24] 96.7 -2.1 -2.7 -5.9
-BAdd [40] 91.0$\pm$0.7 -4.3$\pm$0.4 -1.6$\pm$1.0 -3.9$\pm$0.4
-LfF [31] 97.2 -11.6 -18.4 -63.2
-JTT [26] 95.9 -8.1 -13.3 -40.1
-Debian [23] 98.0 -14.9 -10.5 -69.0
-MAVias 92.8$\pm$0.8 -4.1$\pm$0.6 -2.4$\pm$1.4 -6.7$\pm$1.4
-A.5 Ablation Study
-First, let us report the accuracy of $g_{\boldsymbol{\phi}}(\cdot)$ on the training samples of the urban car class from the UrbanCars dataset, which correlates the target classes with relevant backgrounds and co-occurring objects.
-As shown in Tab.
-13, the $g_{\boldsymbol{\phi}}$ shows a clear distinction in its performance: its accuracy increases from 0% for bias-conflicting samples (i.e., those with a country background and co-occurring object) to 98.42% for bias-aligned samples (i.e., those with an urban background and co-occurring object).
-This result confirms that our tag-based approach produces embeddings that effectively capture bias, providing a foundation for the main model $f_{\boldsymbol{\theta}}(\cdot)$ to focus on learning unbiased features.
-Table 13: MAVias: $g_{\boldsymbol{\theta}}(\cdot)$ accuracy for the UrbanCars training samples belonging to subgroups of urban car class with different biases.
-BG CoObj #samples $g_{\boldsymbol{\theta}}(\cdot)$ acc.
-Country Country 10 00.00
-Country Urban 190 54.21
-Urban Country 190 67.36
-Urban Urban 3610 98.42
-Furthermore, we investigate the impact of the vocabulary size employed for image tagging.
-The original vocabulary RAM uses has a size of 4585 words.
+GroupDro [sagawa2019distributionally] 91.6 10.9 3.6 16.4
+DFR [qiu2023simple] 89.7 10.7 6.9 45.2
+LLE [li2023whac] 96.7 2.1 2.7 5.9
+[…]
+LfF [nam2020LfF] 97.2 11.6 18.4 63.2
+JTT [liu2021just] 95.9 8.1 13.3 40.1
+Debian [li2022discover] 98.0 14.9 10.5 69.0
+[…]
+Figure 6: Impact of hyperparameters $\alpha$ and $\lambda$ on the worst-group accuracy (closed-set protocol) on the Waterbirds dataset.
+5 Ablations & Hyperparameter Tuning
 Tab.
-14 reports the performance of MAVias for different portions of the original vocabulary.
-Notably, using 30% or more of the original vocabulary yields highly effective models, which is due to the dataset’s bias (aquatic or terrestrial environments) that a variety of tags can capture.
-Table 14: Performance of MAVias across varying portions of the original tag vocabulary. Results pertain to the Waterbirds dataset.
-size WG Acc. Avg. Acc.
-10% 78.1 87.1
-20% 79.4 89.0
-30% 91.1 92.0
-40% 93.1 92.3
-50% 93.8 93.2
-60% 93.2 93.5
-70% 92.9 94.2
-80% 93.0 94.3
-90% 93.6 94.5
-100% 93.7 94.5
-Moreover, Tab.
 16 reports the performance of MAVias when computing individual embeddings for each tag and then averaging them compared to the one computing a single embedding for all tags combined.
@@
 Type WG Acc. Avg. Acc.
-separately 93.0$\pm$0.2 93.6$\pm$0.2
-collectively 93.7$\pm$0.4 94.5$\pm$0.4
-Finally, Fig. 7 illustrates the impact of different values of hyperparameters $\alpha$ and $\lambda$ on the MAVias performance.
-Figure 7: Impact of hyperparameters $\alpha$ and $\lambda$ on the Worst-Group accuracy on the Waterbirds dataset.
+[…]
+[…]
+Fig.
+6 illustrates the impact of different values of hyperparameters $\alpha$ and $\lambda$ on the MAVias performance.
+The hyperparameters for all datasets were tuned through the same grid search.
+Overall, optimal $\lambda$ values are close to 0.5 - we noticed that values >0.5>0.5 work better for less biased datasets and <0.5<0.5 for datasets with extreme bias rates.
+$\alpha$ is a much less sensitive hyperparameter, and any value between 0.001-0.1 does not have a severe impact on the results.
+6 Irrelevant Tags
+Tables 17,18,19, and 20 report the lists of irrelevant tags for CelebA, Waterbirds, UrbanCars, and ImageNet9 datasets, respectively.
+Table 17: CelebA irrelevant tags.
+class tags
+non-blonde 3D glasses, Eiffel tower, accordion, adjust, advertisement, alarm clock, album cover, alcohol, amusement park, antler, apple tree, apron, aquarium, archery, arena, arm, army, art exhibition, artist, asparagus, assembly, astronaut, athletic, atm, attach, attend, audi, auditorium, aviator sunglasses, back, backdrop, badge, badminton, balance, ball, ballerina, balloon, bamboo forest, banana tree, banjo, banquet, bar, barbell, barber shop, barrel, baseball, baseball bat, baseball game, baseball glove, baseball hat, baseball pitcher, baseball player, baseball team, baseball uniform, basketball, basketball coach, basketball game, basketball hoop, basketball player, basketball team, bass, bass guitar, bat, bathroom accessory, bathroom sink, batter, battery, beach, beach ball, beach chair, beach volleyball, beaker, beard, beat, bedcover, beer can, beer glass, beige, belly, belt, bench, beret, berry, beverage, bible, bicycle, bicycle helmet, big ben, biker, bill, binder, binocular, birch tree, birthday cake, birthday candle, birthday party, bishop, bite, black, blade, bleacher, blender, blind, blood, blouse, blow, blue, board, boat, bodybuilder, bonnet, book, bookstore, boot, bougainvillea, bouquet, bow, bow tie, bowling ball, boxer, boy, branch, brewery, bride, bridesmaid, bridle, briefcase, brim, broccoli, brooch, bubble gum, building, bull, bullet, bulletin board, bunny, bureau, bus, bush, business suit, businessman, bust, butterfly, button, cabinet, cage, camera, camouflage, can, candle holder, cane, canoe, canopy, cap, cape, car, car door, car seat, card, card game, cardboard box, cardigan, caribbean, carnival, carrot, carry, cash, cat, catcher, cathedral, ceiling, ceiling fan, cello, certificate, chain, chalk, champagne flute, chandelier, chef, cherry, cherry blossom, chessboard, chest, chihuahua, chip, chocolate, chocolate cake, choir, chopstick, christmas, christmas ball, christmas card, christmas hat, christmas tree, church bench, city, city skyline, city street, city view, clasp, classroom, clinic, clipper, cloak, clock, close-up, cloth, clothe, clothing store, cloud, cloudy, club, clutch, clutch bag, coach, cockpit, cocktail dress, coconut tree, coffee, coffee cup, coke, college student, comfort, comic, compete, computer, computer desk, computer monitor, computer screen, concert hall, concrete, condiment, conductor, cone, conference hall, confetti, construction site, construction worker, container, control, convenience store, convention, conversation, convertible, cooler, copy, corn field, couple, course, court, cowboy, cowboy boot, cowboy hat, cream, cricket, cricketer, crochet, crowd, crowded, crown, cruise ship, cube, cup, cut, cylinder, cymbal, dagger, daisy, dark, daybed, debate, debris, decorate, deliver, denim jacket, dentist, desert, detective, diamond, diploma, dirt track, dish washer, diver, dock, doctor, donut, doodle, doorway, dough, dress hat, dress shirt, dressing room, dribble, drink, drive, drone, drum, drummer, drumstick, dry, duchess, dune, dunk, ear, earphone, earth, easter, easter bunny, eclipse, electric outlet, electrician, electronic, emerald, engagement ring, engineer, envelope, equestrian, equipment, evening dress, exercise, exhaust hood, eye, eye mask, eye shadow, eyebrow, face, fair, family photo, fan, fashion model, fashion show, fedora, fence, fern, ferry, fiction, field, fight, figure skater, figurine, file, fill, film premiere, finish line, fish, fishbowl, fishing pole, flag, flannel, flask, flat, flight attendant, flip, floral arrangement, flower girl, flute, fly, foam, foil, food, food processor, football, football coach, football field, football game, football helmet, football match, football player, forehead, fork, frosting, fruit, fruit stand, game, garment, gasmask, gathering, geisha, give, gladiator, glass bowl, glass building, glass jar, glass vase, glasses, glove, goal, goalkeeper, goggles, gold medal, golf cart, golf club, golf course, golfer, gothic, gown, graduate, graduation, graveyard, gray, greyhound, groom, guitar, guitarist, gym, halloween costume, hammer, hammock, hand, handkerchief, hanger, hard rock artist, hardwood floor, hat, hay, head, head coach, headlight, headscarf, hearing, heart, heart shape, hedge, helmet, henna, hillside, hip hop artist, hockey player, hole, honey, hood, hoodie, horse, hospital bed, hospital room, hot, hot dog, hourglass, houseplant, hula hoop, ice cream cone, ice hockey player, icecream, icing, image, infant bed, insect, instrument, interview, ipad, iphone, iron, ivy, jacket, jacuzzi, jar, jaw, jeans, jewel, jewellery, job, jockey, joker, judge, juggle, juice, karaoke, karate, kayak, key, keyboard, khaki, kilt, kimono, kiss, kitchen, kitchen cabinet, kitchen counter, kitchen table, kite, laboratory, lace dress, ladder, lake, lamp post, lanyard, laptop, lash, lasso, laugh, lead singer, leather jacket, ledge, lens, leprechaun, library, license plate, life jacket, lifeguard, light, light switch, lighthouse, lily, limo, liquid, liquor store, list, locker, log, log cabin, logo, love, luggage, machete, machine, machine gun, magnifying glass, magnolia, make, makeup artist, mall, man, mannequin, mantle, mark, martial, martini, martini glass, mask, massage, mat, measuring tape, mechanic, medal, media, messenger bag, metal artist, meter, microphone, microscope, milk, milkshake, mirror, mixer, monarch, money, monk, moped, mother, motorbike, motorboat, motorcycle, motorcycle racer, motorcyclist, mound, mountain, mountain biker, mouse, mouth organ, mouthwash, mp3 player, mud, mug, muscle, music, mustard, napkin, nature, navy, neck, needle, neon, neon light, net, newlywed, news, news conference, nightclub, nose, number, nun, nurse, oar, office, office building, office chair, office supply, office window, officer, open, operate, orange, orange juice, orange tree, out, overcoat, package, paddle, pain, palm tree, panda, panel, paper, paper cup, paper towel, parachute, parasail, parking meter, parrot, party, party hat, pastry, patient, payphone, peach, peak, peel, pen, pencil, pepper, person, pharmacy, phone box, photo, photo frame, picture, picture frame, pigeon, pigtail, pillar, pillow, pilot, pin, pint, pipe, pipe organ, pirate, pitch, pitcher, plaid, plane, plank, plate, play, player, playground, playing card, playingfield, plug hat, podium, point, poker, pole, politician, polo shirt, pomeranian, pop artist, poppy, porch, postcard, powder, power line, prayer, preacher, presentation, president, press room, pretzel, princess, produce, professional, projection screen, projector, prom, propeller, pug, puppet, purple, puzzle, python, queen, race, race car, race track, racket, radiator, radio, raft, rain, raincoat, rapper, razor, read, receive, reception, record player, rectangle, referee, reflection, religious, remote, ride, rifle, rise building, road, robe, rock artist, rock band, rock formation, room, rope, royal, ruffle, rugby player, ruler, run, runner, safety helmet, safety vest, sailor, sake, salon, salute, sand, sandal, sandwich, santa claus, sari, satin, sauna, sax, school bus, scientist, scissors, scoop, scooter, scout, scratch, screen door, scrub, sculptor, sculpture, sea, seashell, seat, seat belt, selfie, service, shadow, shake, shampoo, shape, share, shaver, shaving cream, sheet, shell, shield, shirt, shirtless, shoot, shooting basketball guard, shot glass, shotgun, shoulder, shoulder bag, shower head, sideline, siding, sign, signature, silver medal, sing, singer, sip, sit, skeleton, ski, ski pole, ski slope, skier, skull, sky, sleep, sleepwear, slide, smartphone, smell, snake, snow, snowball, snowman, snowy, soap, soda, softball, soldier, sombrero, sparrow, speaker, speech, speech bubble, spider web, spin, sponge, sports ball, sports coat, spray, square, squirrel, stable, stadium, stage light, stairs, stall, stand, star, stare, steam, stethoscope, stick, stocking, stomach, stool, stop sign, store, strap, straw, straw hat, strawberry, street corner, stretch, string, strip, stripe, student, subway, subway station, subwoofer, suit, sun, sun hat, sunbathe, sunflower field, sunglasses, sunset, sunshade, superhero, surfboard, surfer, surgeon, surgery, suspenders, sweater, sweatshirt, swim, swim cap, swimwear, swing, swinge, switch, swivel chair, syringe, t shirt, t-shirt, table, table tennis, table tennis racket, table tennis table, tablet computer, tackle, take, talk, tan, tape, tassel, tattoo, taxi, tea, team presentation, tear, tee, teething ring, telegraph pole, telescope, television, tennis, tennis match, tennis net, tennis player, tennis racket, test tube, theater, thistle, throne, throw, tie, tiger, tight, tile, tile wall, tissue, tobacco pipe, toddler, toilet paper, tomato, tool, toothpick, torch, totem pole, tower, tower bridge, toy gun, train, train car, train track, train window, tray, treat, tree trunk, trench, tribe, tripod, trophy, trumpet, tub, tulip, turquoise, twig, twist, typewriter, ukulele, underclothes, uniform, vampire, vase, vegetation, vehicle, veil, vessel, video camera, vine, vineyard, violet, violin, violinist, visor, volleyball, volleyball player, waist, wall, wall clock, warehouse, wash, watch, water, water park, waterway, wave, wear, web, wedding, wedding bouquet, wedding dress, wedding ring, wet, wetsuit, whistle, white, white house, whiteboard, wildflower, wind, window sill, windshield, windy, wine cellar, wine glass, winery, winter, witch, witch hat, wood floor, world, wrench, wrestle, wrestler, wristband
+blonde CD, Eiffel tower, Wii, Wii controller, actor, adjust, album, album cover, alley, apron, army, award, baby, backdrop, balustrade, bandage, bandeau, barber, barbie, bartender, baseball bat, bathroom mirror, beach ball, beautiful, beauty salon, bed, bedcover, bikini top, birth, bite, blender, blow, boat, bobfloat, bookshelf, bookstore, bouquet, bow, box, brace, bracelet, braid, bride, bridle, brush, bubble, bust, camera, candy, car seat, cardigan, carpet, casino, cat, catwalk, chihuahua, christmas hat, cigar, city, clip, clock, cloth, clothing store, coach, coat, comb, comfort, convention, cosmetic, crochet, crowd, curl, cut, daisy, dashboard, debate, deliver, denim jacket, diamond, dirt field, doll, dress, dressing room, drive, dye, earring, eat, engagement ring, equestrian, exhaust hood, eyebrow, eyeliner, fan, fashion show, father, fawn, figure skater, flag, flash, flight attendant, flip, floor, flower, folder, freckle, fur, fur coat, game controller, gathering, gift, gift box, girl, graduate, grass, greenery, guitar, headboard, headdress, hearing, hedge, hip, home appliance, hood, horse, hospital bed, hug, hydrant, interview, ivy, jacuzzi, jockey, kangaroo, khaki, kitchen, kitchen counter, knit, lamp, laptop, lash, ledge, leotard, lick, lighthouse, lightning, lipstick, lollipop, lush, makeup artist, microscope, mirror, model, mother, motorcycle, mouth, necklace, neon light, newborn, news, nose, office building, office window, officer, oven, overall, package, palm tree, pastry, pearl, pen, pencil, pet, phone, photo, picture frame, pigtail, pillar, pillow, pilot, pink, podium, poinsettia, pole, ponytail, pool, pop artist, pose, premiere, prom, pug, purple, python, raceway, ramp, razor, relax, ribbon, ring, sailor, salon, savanna, scarf, scooter, screen, seashell, seat belt, selfie, shampoo, shawl, shopping bag, short, shower curtain, slice, smartphone, smell, smile, sparkle, sparrow, speech bubble, sports ball, sports coat, spray, stable, stack, steering wheel, stethoscope, stick, strip, stripe, subway, sunflower, sunshade, suspenders, swim, swimming pool, syringe, tabby, take, tape, tapestry, tiara, toothpaste, toothpick, train window, treat, trench coat, tulip, umbrella, vampire, veil, video camera, violet, violinist, volleyball, volleyball player, wall, water, wedding, wedding dress, wheat field, wheel, wig, wind, wine, wine bottle, woman, wound, wrap, wristband, writing
+Table 18: Waterbirds irrelevant tags.
+class tags
+landbird accordion, airplane window, airship, alcohol, algae, amphitheater, amusement park, anemone, ant, apple, apple tree, aquarium, area, arm, armchair, armor, army, arrow, asparagus, atrium, attach, atv, aurora, author, autumn, autumn forest, autumn leave, autumn park, ax, baby, baby carriage, back, backdrop, backpack, backyard, bag, balance, balcony, ball, balustrade, bamboo, bamboo forest, banana, banana tree, bandeau, bangs, banyan tree, barge, bark, barrel, barrier, baseball bat, baseball hat, basin, bass, bat, bath, beach ball, beach house, beach towel, beam, bear, bear cub, beautiful, beer, beer bottle, belly, bench, bengal tiger, beret, berry, beverage, bicycle, bicycle helmet, biker, bikini, bikini top, bin, birch, birch tree, bird bath, bird cage, bird feeder, blade, blanket, blood, bloom, blossom, blowfish, blue, blue jay, board, boardwalk, boat house, boat ride, bobfloat, bonfire, book, boot, bottle, boulder, bouquet, boy, branch, break, bride, bridesmaid, bronze statue, broom, brown, brunette, brush, bubble, bud, buddha, bug, bundle, buoy, bush, business suit, butterfly, cabana, cabin, cactus, cage, camera, camouflage, camp, campfire, campsite, can, canal, cane, canopy, cape, cardigan, carnival, carp, carpet, carry, cart, carve, catamaran, catch, cave, ceiling, cement, champagne, chase, cherry, cherry blossom, cherry tree, child, christmas hat, christmas light, church, cigar, cigarette, cinema, city park, claw, clear, cloak, cloth, clothesline, cloud, clown fish, cocktail dress, coconut, coconut tree, coffee, coffee cup, collage, color, computer, computer monitor, computer screen, cone, construction site, continent, copy, coral, cosmetic, cotton, cowboy hat, coyote, crack, creek, cross, crowd, crowded, cruise, cub, cup, curb, curtain, curve, cycle, cypress tree, daffodil, daisy, dandelion, dark, dashboard, daughter, debris, decorate, deer, deity, denim jacket, desert, destruction, dew, diaper, dinning table, dirt bike, dirt field, dirt road, dirt track, diving board, doberman, doll, doodle, doorway, dragonfly, drain, drawing, drift, drink, drinking water, drive, dry, duckling, dune, dust, elephant, enclosure, equipment, eucalyptus tree, evening sky, evergreen, face, fairy, fall, family, family photo, fawn, feed, fence, fern, ferris wheel, fiction, field, fighter jet, figurine, finish line, fire, firework, firework display, fish pond, fisherman, fishing net, fjord, flag pole, flame, fleet, flood, floor, flower, flower bed, foam, fog, foggy, foil, folding chair, foot, footprint, footstall, forehead, forest, forest floor, formation, fountain, frog, frost, frosty, fruit tree, fungi, garbage, garden, gas, gazebo, giraffe, glacier, glass building, glass house, glass table, glasses, goldfish, goose, grain, grampus, grass, grasshopper, grassland, grassy, grave, gravel, gravestone, graveyard, gray, graze, green, greenery, greyhound, groom, ground squirrel, grove, guard, guitar, gun, hair, hallway, hammer, hammock, hand, handgun, hang, hay, hedge, helicopter, helmet, herd, hibiscus flower, hide, hike, hiker, hill, hillside, hippo, horse, horseback, hose, house, house exterior, house finch, husky, hut, ice cave, icicle, illuminate, image, individual, insect, instrument, island, ivy, jacket, jaguar, jeans, jeep, jet, jog, jungle, kilt, kiss, kitchen window, kite, koi, ladder, lagoon, lamb, lamp, lamp post, land, lantern, lap, laugh, lavender, lay, lead, lead to, leaf, leash, ledge, lemur, leopard, leotard, lichen, lick, life belt, lifeguard, lift, light, lilac, limb, line, liquid, load, lock, log, log cabin, lotus, lush, magnifying glass, mall, maple, maple leaf, marathon, mark, martini glass, measuring tape, mermaid, miniature, miniskirt, missile, moat, monkey, monocycle, moon, moonlight, morning fog, moss, moth, motorbike, motorcycle, motorcyclist, mound, mountain bike, mountain biker, mountain biking, mountain gorilla, mountain lake, mountain landscape, mountain path, mountain range, mountain stream, mud, mug, mushroom, music, navy, necklace, nectar, net, newlywed, oak tree, office building, officer, oil, open, opera house, orange, orange tree, otter, out, overall, oyster, pagoda, paint brush, palace, palm tree, panda, paper bag, paper lantern, parachute, paradise, parasail, park, park bench, parking lot, passenger train, path, pavement, pavilion, paw, peacock, peak, peanut, pearl, peel, pencil, perch, pergola, pet, petal, pheasant, photographer, pick up, picnic table, pig, pigtail, pillar, pillow, pilot boat, pine cone, pine forest, pink, pipe, pirate ship, pitcher, pitcher plant, plank, plant, plaque, plastic, play, playground, plectrum, plume, pod, pole, pollen, pollution, pond, ponytail, porch, post, postcard, pot, power line, power station, purple, push, python, quail, race, raft, railroad, rain, rainbow, rainbow trout, raincoat, rapid, read, red, red panda, redwood, reed, reel, reflect, reindeer, reptile, resort, restaurant, rice field, rider, rifle, rink, ripple, rise building, river, river boat, robe, rock face, roof, root, round table, rubble, ruler, run, runner, rust, safety helmet, sand castle, sandal, savanna, sax, scooter, scorpion, screen, screen door, sea cave, sea turtle, seashell, seawall, seaweed, sedan, selfie, shawl, shed, sheep, shell, shelter, shine, shipwreck, shopping cart, short, shoulder, shovel, shrine, sign, silverware, sing, sip, sit, skate, skate park, skateboarder, skeleton, skier, skyline, skyscraper, sleep, sleepwear, sleigh, slide, smell, smile, smoke, snail, snake, snorkel, snow, snowboarder, snowman, snowstorm, snowy, spark, sparkle, sparkler, spear, spider web, spike, spire, splash, sprinkler, squirrel monkey, stab, stack, stadium, stand, statue, steam, stem, stingray, stomach, stone building, stool, storm cloud, strap, straw, stream, street light, street sign, string, stroll, structure, stuff, stump, stunt, succulent, suit, sunbathe, sunflower, sunflower field, sunglasses, sunshade, sunshine, surface, surround, suspension bridge, suv, swan boat, sweater, sweatshirt, swim cap, swimming pool, swing, swinge, sword, table, table top, tablecloth, tail, take, tangerine, tank, tape, tattoo, tea, teal, telegraph pole, temperature, temple, tennis racket, tent, terrace, terrain, tie, tiger, tile roof, tire, toe, tool, torch, tortoise, towel, tower bridge, town, track, traffic cone, trail, trailer, trailer truck, train, train car, train track, trainer, travel, tree, tree branch, tree house, tree root, tree trunk, trench, trench coat, tricycle, tripod, trolley, tropic, trout, truck, tub, tube, tugboat, tundra, turquoise, turtle, twig, twilight, umbrella, underclothes, uniform, urban, vegetation, vehicle, veil, volcano, wagon, walk, wall, warning sign, waste, water drop, water feature, water lily, water pipe, water ski, waterfall, waterway, web, wedding, wedding couple, wedding dress, wedding party, weed, wet, wheat field, wheelchair, white house, wild, wildflower, willow, wind chime, wind farm, wind turbine, windmill, window, window sill, wine, wine bottle, wine glass, winter, winter scene, wire, wisteria, wolf, wood, wood floor, wood wall, woodland, world, worm, wrap, writing, yard, yellow, zebra, zoo
+waterbird Border collie, aircraft carrier, album, album cover, algae, amusement park, antler, appear, apple, apple tree, apron, aquarium, army, arrow, aurora, autumn park, ball, ballerina, ballet, ballet dancer, ballet skirt, balustrade, banana tree, barbie, bark, barrier, baseball hat, basket, battleship, beach, beach chair, beach hut, beach towel, beam, beautiful, beaver, belly, bench, berry, bicycle, bikini, bikini top, birch, bird bath, bird feeder, blossom, board, boardwalk, boat, boat house, boat ride, bottle, boy, bridge, broom, brown, brunette, bubble, building, buoy, bus, bus window, bush, cabana, cactus, calm, camera, camp, camper, campsite, can, canoe, car, car mirror, car window, cargo ship, carp, carrot, carry, cave, chain, chair, chase, cheetah, chest, child, chimney, chopstick, cinema, city, city park, city skyline, city view, city wall, cliff, climb, cloth, cloud, cloudy, coast, coat, coconut tree, collage, collect, concept car, cone, corn field, couple, course, crash, crest, crowd, crowded, cruise, cruise ship, curb, cycle, daffodil, dam, dance, dancer, dark, dashboard, daybed, deck, deer, den, denim jacket, destruction, dirt field, dirt road, dirt track, disco ball, diving board, dock, doll, donkey, doorway, driftwood, drive, dry, eel, eucalyptus tree, evergreen, eye, fall, fawn, fence, fern, ferry, field, fighter jet, fin, fishing, fishing pole, fjord, flag, fleet, flip, float, flood, flower bed, fly, fog, foggy, foil, folding chair, food, football field, football game, footprint, forest floor, formation, frog, frosty, frozen, fruit, fur, garden, gazebo, glass table, goal, goggles, goldfish, golf course, goose, grass, grassland, grassy, gravel, gray, graze, guard, gun, habitat, hair, hammock, harbor, hat, hay, hedge, hiker, hillside, hook, horizon, horse, horseback, hose, hotel, house, house exterior, humpback whale, hurdle, hut, hydrangea, ice, ice floe, iceberg, icicle, image, island, jacket, jet, jog, jump, jungle, kayak, kite, knit, koi, ladder, lagoon, lake, lamp post, land, landing, lap, laugh, lay, leaf, ledge, lens, leotard, license plate, lift, light, lighthouse, lightning, log, lotus, machine gun, mallard, mammal, mast, mat, mermaid, mollymawk, moon, moonlight, moose, motorboat, motorcycle, mountain, mountain lake, mud, nectar, needle, nest, night, night sky, night view, oar, office building, open, otter, out, paddle, paling, palm tree, parachute, parasail, park, patch, pavement, peak, pearl, pencil, pepper, perform, person, pet, pheasant, photo, photographer, picnic table, picture, pine cone, pirate ship, pitcher, pitcher plant, plain, plane, plank, plastic, play, point, pole, pollution, pond, pool, poster, power line, power station, prom, puddle, race, raceway, raft, rail, railroad, rainbow, rainbow trout, ramp, ray, rearview mirror, recreational vehicle, redwood, reed, reflect, reflection, relax, resort, ride, rifle, ripple, river bank, riverbed, road, road sign, robe, rock formation, rocky, rodent, rope, rowboat, ruins, run, sail, sailboat, sand, sandal, sari, savanna, schooner, scooter, sea, sea cave, sea lion, seal, selfie, shark, shine, shipwreck, shore, shoreline, silo, sky, skyscraper, sleepwear, smoke, snow, snow leopard, snowy, sphere, splash, spoon, sports car, spray, squat, stand, starling, statue, stingray, stone, stone building, stone house, stool, storm, stormy, straw hat, stream, stretch, strike, string, stump, stunt, sun, sun hat, sunbathe, sunglasses, sunrise, sunset, sunshine, surf, surface, surfboard, surfer, surround, swamp, swan, sweatshirt, swim, swimmer, swimming pool, swing, sword, table, take, takeoff, tank, tape, tarp, taxi, telegraph pole, terrain, thunderstorm, tiara, tortoise, towel, tower, town, track, traffic sign, trailer, train car, train track, trench coat, trim, tripod, tropic, trout, turkey, turquoise, turtle, twilight, urban, vehicle, veil, vessel, video camera, view mirror, vine, warning sign, warship, water, water lily, water ski, waterfall, waterway, wave, weapon, wedding dress, weed, wet, wetland, wetsuit, whale, whale shark, wheat field, white, wig, windshield, winter, winter scene, witch, woodland, zoo
+Table 19: UrbanCars irrelevant tags.
+class tags
+urban-car advertisement, alley, apartment, archway, atm, attach, back, backyard, balcony, balustrade, bear, bin, binocular, bird, black, brick building, brush, building, bus stop, bush, cage, canopy, carpet, catch, ceiling, ceiling fan, charge, chase, chimney, city, city skyline, city street, city view, cliff, cloth, collage, corn field, corridor, country lane, countryside, courthouse, crane, crash, creek, crop, crosswalk, curb, cut, dark, destruction, display, dog house, doodle, door, doorway, dune, dust, enclosure, evergreen, exit, explosion, fence, fill, fire truck, fireplace, flip, flood, floor, fuel, garage, garage door, gas, gas station, gravel, green, greenery, gun, hand, hang, hay, head, headlight, hedge, home, hood, horseback, house, house exterior, hurdle, hut, hydrant, ice skating, illuminate, intersection, jockey, jump, lake, lamp, lamp post, lawn, ledge, lift, lighthouse, line, living room, mat, meter, miniature, mountain, narrow, neon light, night, night view, number, office building, out, outdoor, outhouse, overpass, paling, palm tree, park, parking garage, parking lot, parking meter, parking sign, pavement, peak, pen, pergola, phone box, pillar, pine, pine forest, pink, plant, plug, pole, post, pot, power line, pump, rail, railroad, rain, ramp, ravine, red, reindeer, rink, rise building, road sign, row, sale, sculpture, shed, shelter, sign, signal, silo, skate, skyscraper, smoke, snow, snowboarder, snowy, speed limit sign, stable, stair, stall, stare, stick, stone building, stool, stop sign, store, storefront, street corner, street scene, street sign, stunt, tarp, telescope, terrace, toy, track, traffic light, traffic sign, train, train car, train track, trick, trim, tunnel, umbrella, unicorn, urban, vegetation, vineyard, wall, warehouse, warning sign, water, wheat field, white, white house, yak, yard, yellow, zebra crossing
+country-car advertisement, animal, apartment, arrow, atv, autumn forest, autumn tree, back, balustrade, bamboo forest, barrier, bat, bear, bed, bend, bicycle, black, boardwalk, boulder, bride, bridge, brown, brown bear, buffalo, build, bull, bus stop, bush, cage, calf, cart, cattle, chase, chimney, city street, cliff, climb, cloud, cloudy, coach, construction site, corridor, country lane, couple, cow, cowboy, crane, crop, cross, curve, daffodil, dam, deer, desert, desert road, dirt bike, dirt field, dirt road, dirt track, dog, donkey, door, dust, elephant, enclosure, evergreen, farmland, fence, field, fire escape, fire hose, fire truck, flag, flower, forest, forest path, forest road, gas station, gasmask, goat, goggles, grass, grassy, grave, gravestone, graveyard, gray, graze, green, gun, hallway, handstand, head, hedge, highway, hill, hillside, hook, horse, horse cart, horseback, hut, ice cream cone, jump, lamb, land, lay, leash, lift, lighthouse, line, log, lush, milk, minibus, moose, mound, mountain, mountain path, mud, mural, night sky, out, parachute, parking, path, peak, pig, pine, pine forest, pink, plain, plectrum, plow, pony, pull, pump, railroad bridge, ram, redwood, reflection, relax, rhinoceros, rice field, ride, rise building, road, road sign, rock formation, rocky, rural, sand, savanna, scaffold, schooner, sculpture, sea, sheep, shelter, signal, silver, sit, sky, snow, snowy, speed limit, speed limit sign, squirrel, stable, stair, stairwell, stand, star, stare, stone, stool, storefront, stormy, street corner, street light, stretch, stuff, sun, sunset, swing, telegraph pole, terrain, toy, track, traffic cone, trailer, travel, tree, trench, trick, trumpet, umbrella, unicorn, view mirror, vineyard, wagon, walk, warehouse, warning sign, water tank, water tower, white, wood, yak
+Table 20: ImageNet9 irrelevant tags.
+class tags
+Dog CD, French bulldog, antler, appliance, area, arm, armchair, back, backdrop, backyard, bag, balance, barn, barrel, barricade, barrier, baseball hat, basin, basket, bath, bathroom, bathroom sink, bead, bed, bedcover, belt, bench, beret, bib, black, blanket, blood, blue, boardwalk, boat, bookcase, bookshelf, boulder, bow tie, bowl, boxer, boy, break, broom, brown, brush, bubble, building, bureau, bush, cabinet, cap, car, car seat, car window, carpet, carrot, carry, cat, cat food, cement, chair, child, christmas, christmas hat, cloth, cocktail dress, comfort, computer, computer chair, computer desk, computer monitor, concrete, cone, contest, corridor, counter top, course, court, crack, cube, curb, dachshund, daisy, dandelion, deck, dirt field, dirt road, dirt track, dish washer, dock, doll, door, doorway, dot, drawer, dress, dresser, drink, duffel, eat, enclosure, eye, face, fairy, farm, feed, fence, fern, finish line, fireplace, flower, flower bed, folding chair, football, footprint, forest, fridge, frisbee, frost, frosty, garden, gift, girl, glass door, goggles, gravel, green, greyhound, guinea pig, hair, hallway, hand, handkerchief, hardwood, hardwood floor, hassock, hay, head, headscarf, hide, hillside, home appliance, hose, house exterior, hug, hurdle, jeans, judge, kiss, kitchen, kitchen cabinet, kitchen counter, kitchen floor, kitchen island, kitchen sink, lamb, lap, laundry basket, lawn, leather jacket, ledge, living room, log, lollipop, lush, magnet, man, measuring tape, milk, miniature, moose, motorbike, motorboat, motorcycle, motorcyclist, mound, neck, necklace, obstacle course, office, office chair, office supply, open, out, paling, paper towel, park, pasture, patch, path, pavement, pen, pencil, person, photo frame, picnic table, picture frame, plant, plate, pointer, pole, polka dot, poodle, porch, pot, quilt, racket, rail, ramp, red, reindeer, river, road, rocking chair, rocky, roll, room, round table, rural, scarf, screen door, seat, seat belt, shadow, shake, shirt, siding, sign, sink, ski, ski pole, sky, sleeping bag, sleepwear, slide, smile, snow, snowy, soap, sports ball, spray, stand, stare, stethoscope, stick, stone, stool, strap, stuff, stump, sun, sunglasses, swab, sweater, sweatshirt, tabby, table, take, tan, teddy, tennis, tennis court, tie, tile, tool, towel, trainer, trumpet, tub, tube, tulip, tusk, umbrella, vegetation, wagon, wash, watering can, wear, wet, white, white house, window, window sill, wine glass, winter, woman, wood, wood floor, wood wall, wool, wrap, wrapping paper, wrestle, yellow
+Fish algae, appear, aquarium, attach, bag, bait, ball, banana, baseball hat, basin, bass, bat, bath, bathroom, bathroom accessory, bathroom sink, beach, beach chair, bean, beard, beautiful, beetle, billiard table, bin, bite, black, blade, blood, blue, boat, bowl, boy, brush, bubble, bug, buoy, cactus, cake, calm, camouflage, camp, campfire, campsite, canoe, cat, catch, caterpillar, cave, ceiling, chair, child, chocolate cake, clear, cliff, cloth, coast, color, comb, container, coral, coral reef, couple, coverall, cowboy hat, crack, creek, dark, display, dive, diver, dock, doll, dolphin, dress, dry, dye, enclosure, entertainment center, eye, fang, fedora, fin, fish market, fisherman, fishing net, fishing pole, flat, float, folding chair, foot, footprint, frosting, frying pan, gasmask, gecko, girl, glass bowl, glass table, grampus, gray, green, hairbrush, hand, hat, head, hole, humpback whale, ice, image, jacket, joker, jumpsuit, kayak, koi, lagoon, lake, land, laugh, ledge, lid, life jacket, lush, machete, man, market, mask, mine, mound, mouse, mud, muscle, net, night, overall, package, pad, paddle, pancake, peak, pearl, peel, penguin, pepper, person, pet, pharmacy, phone, photo, pink, plant, plastic, plate, platter, pod, pole, polka dot, polo shirt, pond, pool, pose, profile, puddle, purple, raincoat, ram, reel, river, rock formation, rocky, sack, safety vest, sale, sandal, sea, sea cave, sea lion, sea urchin, seabed, seal, seashell, shine, shirt, shirtless, shore, shoreline, showcase, silver, sit, skate, smile, snorkel, snow, spike, sponge, squat, squid, stab, stand, stem, stingray, stone, stool, strainer, strap, stream, stuff, sun, sun hat, sunglasses, surface, suspenders, sweatshirt, swim, swimming pool, swimwear, swordfish, table, tail, tank, tape, tarp, teddy, tent, toe, toilet bowl, toilet seat, tool, torch, tortoise, toy, tree, tripod, tropic, trumpet, tube, turquoise, turtle, tusk, twig, underwater, urinal, vegetation, vessel, vest, water, waterway, wear, weed, wetsuit, whale, whale shark, white, woman, wood, yellow
+Bird area, attach, baby, balustrade, bark, beach, bean, bed, bend, berry, bill, birch tree, bird bath, bird cage, bird feeder, bird house, bird nest, birdbath, blanket, bloom, blossom, blue, boulder, branch, brown, brush, bud, building, bunk bed, bureau, bush, cabinet, calm, car mirror, car window, catch, ceiling, chase, cherry blossom, cherry tree, circle, clear, climb, close-up, cloud, coast, comb, cone, counter top, couple, creek, crest, cross, debris, dirt field, dock, driftwood, drink, dry, enclosure, evergreen, eye, face, feed, fence, field, fir tree, fish, flap, flight, float, flood, flower, flower bed, fly, food, foot, forehead, forest, garbage, grain, grassland, grassy, gravel, gray, graze, greenery, gull, hand, hay, head, hide, hillside, house finch, kitchen counter, kite, ladder, lake, land, landing, lap, laugh, lavender, ledge, leg, lizard, metal, mollymawk, mound, mouth, mud, mulch, neck, nectar, nest, night, nose, nut, open, osprey, ostrich, out, paling, paper, patch, peak, peanut, pen, penguin, perch, pillar, pillow, pine, pine cone, pink, plain, plank, plant, plow, portrait, post, power line, puddle, purple, raft, rapid, rearview mirror, red, reflection, relax, rice field, ripple, river, rock formation, rocky, roof, room, rope, sand, savanna, sea, seabird, seal, seaweed, seed, shirt, shore, shoreline, sip, sit, sky, smile, snowy, spread, spruce, stem, stick, stilt, stone, stretch, string, sunflower seed, swim, table, tail, teal, telegraph pole, thistle, tile, tile roof, toy, tree, tree branch, tropic, turkey, twig, vegetation, view mirror, walk, water, water lily, wave, weed, wet, wetland, wildflower, window sill, windshield, wing, wire, wood, yard, yellow, zoo
+Vehicle advertisement, air field, alcohol, alley, ambulance, amusement ride, animal, antique, approach, arena, armor, auto show, baby, back, balance, barn, barrel, barricade, barrier, basin, beach, beach chair, beautiful, bench, bend, bicycle helmet, biker, bin, bird cage, black, blow, blue, boat, boot, bottle, box, boy, brick building, bridge, broom, brown, brunette, buffalo, building, bull, bullet, bumper, cabin, cabin car, cage, camper, campsite, cane, canopy, car show, cardboard box, carpet, cat, catch, cattle, ceiling, cement, chair, check, child, city, city square, city street, classic, cliff, cloud, cloudy, coast, cock, commuter, cone, construction site, construction worker, container, cooler, countryside, coupe, couple, course, cow, cowboy boot, cowboy hat, crane, crate, crop, cross, crowd, crowded, curb, curve, cycle, cylinder, dark, dashboard, decker bus, decorate, desert, desert road, destruction, dirt bike, dirt field, display, dock, dog house, donkey, doodle, door, doormat, doorway, dozer, dress, drift, driveway, dust, elephant, emergency, emergency service, enclosure, envelope, equipment, evergreen, excavator, face, fence, field, figurine, fill, fire department, fire hose, fire station, fire truck, fireman, flag, flame, flip, flood, floor mat, flower, fly, food truck, forest, garage, garage door, garbage, garbage truck, garden, garden hose, generator, girl, go, goat, goggles, gold, golf cart, grass, grassland, grassy, gravel, gray, green, handcart, hangar, hat, hay, headlight, heart, hedge, helmet, hide, highway, hill, hillside, hip, hood, hose, house, house exterior, hut, intersection, jacket, job, jockey, jump, jumpsuit, karting, kayak, ladder, lake, lamp, lamp post, land, lap, lawn, lay, lean, leash, lego, lifeguard, lift, lighthouse, livestock, load, loading dock, lock, lorry, machine, magazine, mall, mat, miniature, miniskirt, missile, model, model car, monocycle, monster truck, motorcycle racer, motorcyclist, mound, mountain, mountain path, mower, mud, mulch, mule, muscle car, museum, narrow, night view, number, obstacle course, office building, open, operate, overall, overpass, palette, palm tree, panel, parade, paramedic, park bench, parking, parking lot, pass, passenger, passenger train, patch, pavement, peak, pen, pet, photo, pick up, picnic table, pillow, pine, pink, pipe, plain, plant, plaque, platform, plaza, plectrum, plow, plume, pole, police, police station, pose, pot, power line, puff, purple, race, race track, rail, railcar, railroad, railroad bridge, railway station, rain, rake, ramp, red, repair, rickshaw, rifle, river, road, rock face, rock formation, rocket, room, rust, saddle, safari, safety vest, sand, savanna, scaffold, scale model, schooner, sea, seawall, security, shed, shelf, shipping container, shirt, shoreline, short, shovel, showcase, sidecar, sign, silver, ski equipment, ski slope, skier, sky, sleep, sleigh, slope, smile, smoke, snail, snow, snowboard, snowboarder, snowmobile, snowplow, snowstorm, snowy, sock, solar battery, spectator, speedboat, spin, spray, station, steam, steam engine, steam train, stilt, stocking, stone building, stormy, strap, straw, straw hat, street artist, street corner, street scene, street sign, stretch, string, stunt, sun hat, sunset, swimwear, table, tablecloth, tape, tarmac, tarp, teach, teal, telegraph pole, terminal, tile wall, tire, tool, toolbox, tour bus, toy, toy car, track, tractor, traffic cone, traffic light, trail, train track, travel, tree, tree trunk, trench, trick, trolley, troop, trumpet, tub, turn, turquoise, umbrella, underclothes, urban, vegetation, vespa, vessel, vest, wagon wheel, walk, wall, warehouse, waste, water, water tank, watering can, weapon, wear, weed, wet, wheel, wheelie, wildflower, wind farm, wind turbine, windshield, wine bottle, winter, wiper, wood, wood wall, worker, wreath, writing, yard
+Reptile algae, amusement park, aquarium, arm, armadillo, baby elephant, back, bag, balloon, banana, banana tree, bank, bark, barrel, baseball glove, baseball hat, basin, bath, bathroom sink, beach, beak, bin, bird, black, blanket, blowfish, blue, bottle, boulder, box, bracelet, branch, break, brick, bronze statue, brown, bubble, bug, building, bull, bush, cage, car, car window, cardboard box, carnival, carp, cat, ceiling, cement, chair, chase, cigarette, claw, clear, climb, cloth, clown fish, coast, coconut tree, coin, computer, computer desk, concrete, container, coral, coral reef, crack, creek, crochet, cross, curl, dart, dashboard, desert, dinosaur, dip, dirt field, dirt road, dirt track, display, diver, draw, driftwood, dry, duckling, dune, eagle, enclosure, eye, fang, fern, field, figurine, fish, float, floor, flower, flower bed, food, footprint, forest floor, fossil, fruit, garden, glass table, goggles, grass, grasshopper, grassy, gravel, green, greenery, habitat, hamster, hang, hardwood floor, hay, head, hedge, hide, hill country, hippo, hood, hunt, hut, jaw, key, keyboard, knot, land, lap, lay, leaf, leash, ledge, lego, lid, limb, log, lush, mammal, man, marine, mat, measuring tape, meat, melon, miniature, moose, moss, moth, mouse, mud, mulch, museum, neck, nest, net, night, office supply, open, orange tree, package, pad, palm tree, paper towel, parrot, patch, path, pavement, paw, peak, pebble, pen, penguin, penny, perch, person, petal, photo, pillar, pillow, pink, pipe, plank, plant, plastic, play, pole, polka dot, pond, pool, puddle, purple, rail, ram, rat, red, reed, reef, reflection, relax, rhinoceros, ride, ripple, river, river bank, road, rock face, rock formation, rocky, rodent, sand, sandal, school, screen door, sculpture, sea, sea lion, seal, seashell, seaweed, shadow, shape, shell, shoe, shore, shoreline, showcase, shrub, sink, sip, sit, skeleton, skull, sky, slider, slipper, slug, snail, snow, spiral, spray, stand, stare, statue, stone, stool, strap, straw, stuff, stump, sun, sun hat, sunbathe, surf, surface, swamp, swim, swimmer, swimming pool, table, tank, terrarium, toilet bowl, toilet paper, toilet seat, toy, tray, tree, tree branch, tree trunk, trumpet, tub, tube, twig, underwater, unicorn, up, vegetable, vegetation, walk, wall, wash, watch, water, wave, weed, wetsuit, whale, white, window sill, windshield, wood, wood floor, world, wrap, wristband, yellow, zoo
+Carnivore CD, appear, apple, apple tree, aquarium, area, arm, backdrop, bag, balance, ball, bamboo, bamboo forest, bank, basket, bath, bathroom accessory, bathroom sink, beach, beak, bear cub, beaver, belly, bin, birch tree, bird cage, bite, blanket, blood, boat, bottle, boulder, break, brush, bureau, bush, cabinet, cage, can, car, car window, catamount, catch, cave, chair, chase, chopstick, cigarette, claw, close-up, closet, cloth, container, counter top, cowboy hat, crane, crate, creek, cub, cube, curl, daisy, dandelion, dark, deer, den, dirt field, dirt road, dirt track, dog breed, dog house, doorway, dot, dresser, drink, dry, eat, enclosure, evergreen, eye, fall, fang, faucet, fence, field, fir tree, fish, flower, flower bed, foot, footstall, forest, frosty, fruit, fur, grass, grassland, grassy, graze, green, greenery, greet, ground squirrel, guard, habitat, hamster, hand, hay, hide, hill, hillside, hole, ice floe, iceberg, illuminate, image, jaw, jump, jungle, lake, lamb, land, lap, laugh, lay, ledge, lick, log, lush, mammal, man, mane, meat, mound, mountain stream, mouth, mud, mulch, nest, night, night view, open, paper bag, path, pavement, peak, pen, pencil, perch, photo, pine, pink, plain, plant, plastic, play, pole, pond, pool, portrait, post, postcard, profile, raft, rail, ram, red, red panda, reed, relax, river, river bank, riverbed, road, rock face, rock formation, rocky, roof, run, rural, savanna, screen door, sea, sea lion, seal, shadow, sheep, shelter, shirt, sink, sky, sleep, smile, snow, snowball, snowy, spruce, squat, squirrel, stand, stem, stick, stone, straw, stream, street dog, stretch, stripe, stuff, stump, sun, swamp, swim, tabby, tank, terrain, tiger, tire, toe, tree, tree trunk, trout, trumpet, tub, tube, tundra, twig, vegetation, walk, wall, waterway, weed, wild, wildflower, window sill, winter, woman, wood, wrestle, yard, yellow, zoo
+Insect acorn, alcohol, apple, apple tree, asphalt road, attach, autumn leave, ball, banana leaf, banana tree, bark, bat, bathroom, bathroom sink, beautiful, berry, bird, bird nest, bite, black, blackberry, bloom, blossom, blue, boulder, break, brick, broom, brown, brush, bumblebee, bush, butterfly house, cactus, candy cane, cane, car, cardboard, cardboard box, carpet, cat, ceiling, cement, chestnut, climb, close-up, cloth, cloudy, color, concrete, container, couple, crack, cricket, curb, daisy, dandelion, dirt field, dot, dry, evergreen, eye, feather, fence, fern, field, figurine, flap, floor, foot, forest floor, frog, fruit, fruit tree, garden, gasmask, glass bead, glass jar, glass vase, goggles, gold, grain, grass, gravel, gray, green, greenery, hand, hang, head, hide, hornet, hummingbird, image, ivy, jar, jump, land, larva, lavender, leaf, ledge, lemon, lid, lilac, limb, lush, maple leaf, marble, mat, measuring tape, microscope, mud, mulch, napkin, needle, nest, night, nose, nut, orange, orange tree, pad, paint brush, paper towel, parrot, patch, pavement, peel, peony, perch, person, pest, pet, petal, photo, pillar, pillow, pine, pink, plant, plastic, pole, polka dot, pond, poppy, post, pot, puddle, rail, red, reed, roach, road, rock face, rocky, roll, rope, rose, ruler, sand, scale, screen door, sea, seed, shadow, sign, sit, skin, sky, slug, spider web, spruce, stab, stand, stare, stem, stick, stone, stool, string, stripe, stump, swab, swallowtail butterfly, swirl, tabby, tank, thistle, tile wall, tissue, toilet paper, toilet seat, tool, tree, tree branch, tree frog, tree trunk, vegetation, violet, walk, wall, water, water drop, wave, web, white, wildflower, window, wire, wood wall, writing, yellow
+Instrument CD, Eiffel tower, alcohol, alley, altar, amplifier, animal, antique, arch, armchair, army, art, artifact, assembly line, autumn leave, baby, backdrop, bag, balcony, ball, balustrade, band, bangs, banjo, bar stool, barrel, baseball bat, baseball glove, baseball hat, basement, bassist, bat, bath, bathroom, bathroom accessory, bathroom sink, battery, beach, bead, beam, beard, bedroom, beer, beer can, belt, beret, beverage, billard, billiard table, bird, birthday, birthday cake, bite, blade, blanket, blender, blowfish, boardwalk, boat, boiler, book, bookshelf, boot, bottle, bottle screw, bouquet, bow, bow tie, bowling ball, braid, brass, brick, brick building, brooch, broom, brown, brunette, brush, bubble, bullet, bundle, bureau, bust, cabinet, cage, cake, camp, can, cane, canopy, cape, carnival, carpet, carrot, carry, carve, cat, cathedral, ceiling, ceiling fan, chain, chandelier, chapel, check, chestnut, chicken, chocolate cake, choir, chopstick, christmas hat, church bench, cigar, cigarette, clay, cloak, close-up, closet, cloth, cloudy, club, clutch, clutch bag, coast, cock, cocktail dress, comb, computer, concert, concert hall, conductor, connect, container, copper, corner, cosmetic, cosplay, costume, cotton candy, couch, counter top, cowboy hat, crowd, crowded, crumb, cue, cup, curb, curtain, cylinder, dam, dance, dark, dart, daylight, decorate, design, dirt field, disco ball, display, doll, door, door handle, draw, drawer, dress, dress hat, dress shirt, dresser, drink, dumbbell, eagle, earphone, earring, elastic band, electronic, equipment, face, fair, fan, faucet, feather, fedora, festival, field, figurine, flag, flap, float, flower, folding chair, fondant, foot, football field, football game, footrest, footstall, fountain pen, frosting, fruit, frying pan, garment, gathering, geisha, glass bead, glass bowl, glass vase, glass window, glasses, goggles, gold, grass, grassy, gray, green, hair, hall, halloween, hand, handle, hang, hardwood, hardwood floor, harpsichord, hassock, hat, hay, head, headdress, headscarf, heart, hen, hill, hillside, hole, home appliance, hook, hose, hospital bed, icing, image, iron, ivory, jacket, jeans, joker, juice, keyboard, khaki, kilt, kit, kitchen counter, kitchen utensil, kitchenware, knot, lamp, lamp shade, lap, laptop, laptop keyboard, laugh, lead singer, lean, lecture hall, lick, lid, light, liquid, living room, load, log, loom, lush, machine, magnet, magnifying glass, make, mallet, maraca, march, marching band, match, measuring tape, mechanical fan, megaphone, milkshake, miniature, mirror, mixing bowl, mixture, model, molding, money, monk, mother, mouse, mouth, muscle, museum, music stool, music studio, musician, napkin, navy, neck, necklace, nose, nun, nut, officer, onion, open, outdoor, overall, overpass, owl, package, pad, paddle, palace, palette, pan, pant, paper towel, parade, park, parrot, pearl, pen, pencil, pendant, pendulum, perch, perform, phone, photo frame, picnic table, picture frame, piggy bank, pigtail, pillar, pin, pink, pipe, pirate, plaid, plank, plant, plaque, plastic, plate, platter, play, plaza, plug, plum, pocket watch, point, poker table, ponytail, pool, porch, pose, poster, pot, preacher, pulpit, purple, pyramid, quilt, racket, radiator, razor, razor blade, recording studio, reed, reel, reflection, relief, rifle, road, robe, rock band, rock concert, rocking chair, roller skate, rope, row, rust, sailor, sand, sauna, saw, scarf, scoop, screw, screwdriver, scroll, sculpture, sea, seashell, selfie, shape, shawl, sheet, sheet music, shelf, shell, shelve, shield, shirtless, shore, shoreline, short, shoulder, showcase, shrine, side, signature, silver, sing, singer, sink, skate, skateboard, skateboarder, skeleton, skewer, ski pole, skier, sleep, sleepwear, smartphone, smile, smoke, snowboarder, soda, sombrero, sound, spatula, speaker, spear, sphere, spin, spoon, sport equipment, stack, stage, stage light, stair, stapler, stem, stick, stir, stone, stone building, strap, straw, straw hat, strawman, street artist, street corner, street vendor, strip, student, stump, subway, subway station, suit, sun hat, sunglasses, surface, surfboard, sweater, sweatshirt, swimming pool, swing, swinge, t-shirt, table, table lamp, tablecloth, tassel, tea pot, teal, telescope, tennis racket, tent, theater, tile wall, tin, tire, toddler, toilet paper, toilet seat, tool, toothbrush, toothpaste, toothpick, traditional, transporter, tray, tree, tribe, trick, tripod, trophy, tub, tube, turquoise, tusk, twig, uniform, utensil, velvet, vendor, vest, violinist, walk, wall, wall clock, wall lamp, warehouse, water, wig, wind chime, window, wine bottle, wine glass, wing, wire, witch, wok, woman, wood, wood floor, wood wall, wooden spoon, workbench, workshop, wristband, writing, yellow
+Primate appear, apple, attach, baby, baby elephant, back, balance, ball, balustrade, banana tree, bandage, barrel, basin, basket, bat, bath, bench, berry, bird cage, bite, black, blade, blanket, blue, boardwalk, bookshelf, bottle, boulder, branch, break, broccoli, brown, brown bear, building, canopy, car, car window, card, carrot, cat, catch, cement, chair, chase, clarinet, claw, cliff, close-up, cob, coconut, computer, container, couple, cub, curb, dinning table, dirt road, dirt track, display, dog, dog house, drink, dry, eat, enclosure, eye, face, fang, field, flower, flute, fruit, fur, grassland, greet, greeting card, hair, hammock, hand, hang, hat, hay, head, hedge, hook, hot, howler, huddle, hug, husky, hyaena, impala, jump, karate, koala, land, lap, laptop, laugh, lay, leash, ledge, leprechaun, lick, log, mammal, man, mane, martial, milk, motorbike, motorcycle, motorcyclist, mound, mountain, mouth, neck, neckband, night, nose, open, owl, palm tree, panda, park bench, parrot, patch, peak, peel, perch, person, pet, picnic table, pillar, plain, plank, plant, play, pole, pool, poppy, power line, profile, rail, ramp, red, red panda, relax, ride, rock face, rock formation, roll, rope, rope bridge, sale, sand, savanna, scooter, sea, shadow, shirtless, shoulder, showcase, siamese, sit, sky, sleep, sloth, smile, squat, stab, stare, steam, stem, stick, stone, stool, strap, straw, stretch, string, stuff, stump, sun hat, swing, swinge, table, tablecloth, tassel, telegraph pole, tie, tile wall, toy, treat, tree, tree branch, tree trunk, trumpet, tub, tusk, twig, wear, weed, white, window, window sill, wire, wood, yellow